### v1

# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

Abstract
Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at different rates. However, in such formulation, the point-wise query of field values does not naturally match the point spread function (PSF) of a given pixel. In this work we present a novel way to design neural fields such that points can be queried with a Gaussian PSF, which serves as anti-aliasing when moving across resolutions for ASSR. We achieve this using a novel activation function derived from Fourier theory and the heat equation. This comes at no additional cost: querying a point with a Gaussian PSF in our framework does not affect computational cost, unlike filtering in the image domain. Coupled with a hypernetwork, our method not only provides theoretically guaranteed anti-aliasing, but also sets a new bar for ASSR while also being more parameter-efficient than previous methods.
1 Introduction
For years, learning-based image super-resolution (SR) methods have achieved ever better results. But unlike interpolation methods, which allow us to resample a given image at any resolution, such methods require retraining for each desired scaling factor. In recent years, several supervised methods for arbitrary-scale SR (ASSR) have been proposed to allow end-users to choose any scaling factor, rendering such systems much more flexible [17]. Notably, Chen et al. [11] pioneered the use of neural fields for single-image super-resolution (SR) with LIIF, exploiting the continuous representation to enable SR at arbitrary scaling factors. Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [43]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and analytically differentiable. Among others, they have been widely adopted for 3D reconstruction with radiance [30], occupancy [29, 35], and distance fields [33, 37].
Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components. We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies. This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
Neural fields offer a compelling representation of continuous signals. However, for many tasks it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3]. This is not trivial [3, 4, 16, 5]. For ASSR, where the signal should be observed at various scales and is prone to aliasing, a neighborhood-dependent observation model is desirable. Using appropriate observation modeling has not only the potential to reduce aliasing but also to more accurately reflect the physical processes of image acquisition. For this reason, Chen et al. [11] already took a first step towards learning multiscale representations through cell encoding. For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the location $\mathbf{x}$ as a tool to control different frequencies in $\Phi$.
In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction. Furthermore, this filtering incurs no additional overhead: The cost of querying a field’s value does not depend on the size of the anti-aliasing filter. The key insight is that when using sinusoidal activation functions [38], one can decay different components at different rates based on their frequencies, according to Fourier theory. Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, an ASSR method with neural fields that uses a novel thermal activation function (see Fig. 1). This formulation makes it possible to analytically compute Gaussian-blurred versions of a field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius. This is equivalent to querying $\Phi$ with a Gaussian PSF. Thera outperforms competing methods in most of our experiments, is far more parameter-efficient, and provides theoretical guarantees w.r.t. anti-aliasing and multi-scale representation. We demonstrate that a lightweight variant of our algorithm still outperforms most previous methods while using 2-4 orders of magnitude fewer parameters. To provide theoretical guarantees for our thermal activations, certain constraints are imposed on the field’s neural architecture. Fortunately, these constraints can easily be overcome with the help of a hypernetwork [14], so that our method can seamlessly integrate a wide range of well-proven SR backbones. In summary, our main contributions are:
  1. 1.
    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and as a consequence anti-aliasing with minimal overhead.
  2. 2.
    We use thermal activations to build Thera, a novel method for ASSR which is superior to previous methods while being more parameter efficient and providing theoretical guarantees w.r.t. multiscale representation.
2 Related work
Arbitrary-scale SR. In recent years, research on arbitrary-scale super-resolution (ASSR) methods has gained attention. In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork, conditioned on the desired scaling factor. An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by a neural field that is locally conditioned with features extracted from the low-resolution image. The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution. Following work has since built upon the LIIF framework. For instance, UltraSR [45] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields. LTE [22] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [41] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. Most recently, methods have integrated attention mechanisms into their implicit representation, enabling the neural field to adaptively attend to image features depending on the query coordinates and scale factor, improving reconstruction quality [8, 10]. These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data. In this paper, we explore a way to directly integrate such a model into the neural field representation.
Implicit neural representations. A neural field is a neural network trained to map real-valued coordinates onto a field’s values. Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 38, 39, 11, 22, 12], 3D scenes (e.g., represented as signed distance fields [33, 38, 37, 42], occupancy fields [29, 35], LiDAR fields [19], view-dependent radiance fields [30, 3, 4, 5]), or digital humans [46, 49, 9, 44, 13]. We refer the interested reader to Xie et al. [43] for a comprehensive survey of neural fields in vision and beyond. Frequently, it is desirable to impose some prior over the space of learnable representations. A common approach for such conditioning of neural fields is encoder-based inference [43], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 41, 8, 10]. The encoded latent variables $\bm{z}$ are then used to condition field parameters $\bm{\theta}$, for instance by simply concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork  [14]. The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, mapping latent codes to field parameters. An early example of this is Sitzmann et al. [38], and this approach is gaining popularity [43].
Anti-aliasing in neural fields. Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [3] proposed integrating the positional encoding with Gaussian weights, which reduced aliasing in NeRF. Improvements were later proposed for unbounded scenes [4] and to improve efficiency [16]. Barron et al. [5] tackle anti-aliasing for the increasingly popular Instant-NGP [31] approach. In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
3 Method
To understand how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis. While we stick to the 2D case here for simplicity, all theory is valid for signals with any number of dimensions. We refer the reader to [32] for an in-depth analysis of sampling and aliasing.
To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ in 2D image space, $\mathbf{x}\in\mathbb{R}^{2}$, it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate. One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, and then sample the band-limited signal $f\circledast g(\mathbf{x})$. Here, we show that one can naturally integrate this behavior into a neural field’s architecture. The key to this is knowing that if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor. This equivalence is depicted in Fig. 2. While decomposing a given signal into sinusoidal components is normally a costly operation, we show that we can avoid it by designing a neural field in which the decomposition itself is the native representation of the signal.
3.1 Thermal activation functions
The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation. Other implicit representation methods, such as SIREN [38], optimize the parameters $\bm{\theta}_{\Phi}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$. Our key insight is the following: What if we take this field to represent the temperature distribution on a surface at a given time $t_{0}$? It is well known how to solve the heat equation forwards in time – its solution (on an isotropic surface) is given by a Gaussian blur of the state at $t_{0}$, where the blur radius depends on the time difference $t-t_{0}$. Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
$\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;,$ (1)
where $\kappa>0$ is the thermal diffusivity of the medium. The solution for $0\leq t<t_{0}$ is, however, ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
Figure 2: If we have access to the sinusoidal decomposition of a signal, filtering can be done analytically by scaling each component appropriately. We design our neural field so that we have access to such a decomposition at no extra cost.
A naive attempt to ensure that $\Phi$ respects the heat equation in Eq. 1 would be to include an additional loss term: $\left|\frac{\partial\Phi}{\partial t}-\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi\right|$. Despite our best efforts, this approach was unsuccessful. Nevertheless, it turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
$\mathbf{y}=\mathbf{W}_{2}\cdot\xi(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t)+\textbf{b}_{2},$ (2)
where $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$, described in more detail later; and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time. We set $t=0$ to be the (unobservable) initial thermal state of a given system represented as a sum of sinusoidal components:
$\xi(\mathbf{z},\nu,\kappa,0)=\sin(\mathbf{z}).$ (3)
For $t>0$, each component is decayed according to its wave number, the thermal diffusivity constant $\kappa$, and $t$:
$\xi(\mathbf{z},\nu,\kappa,t)=\sin(\mathbf{z})\cdot\exp(-\left|\nu\right|^{2}\kappa t).$ (4)
To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally over time, regardless of their orientation). Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
There is an ideal bijection between the sampling rate $f_{s}$ and $t$. At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$). A low-pass filtered version of the signal is observed for $t>0$. To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$. The relationship between the cut-off frequency of the filter and $t$ is controlled by $\kappa$, which we can freely set to any positive number. For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$. Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
$\kappa=\frac{\sqrt{\log(4)}}{2f_{s}\pi^{2}}.$ (5)
To subsample the signal $\mathcal{D}$ by a factor $S$, $\Phi$ should be sampled at
$t=S^{2}.$ (6)
The derivations of these values can be found in Appendix A.
In our formulation, $\mathbf{W}_{1}$ can be seen as a frequency bank whose components (“basis functions”) make up the signal $\Phi(\mathbf{x},0)$, and the vector $\mathbf{b}_{1}$ represents the phase shifts of those components. The matrix $\mathbf{W}_{2}$, with one row per output channel, contains the initial ($t=0$) magnitudes of these components, and $\mathbf{b}_{2}$ is the global bias of $\Phi$ per channel.
Image domain
Refer to caption Refer to caption Refer to caption
Original Anti-aliasing for $\times$1/4 Anti-aliasing for $\times$1/16
Thera
Refer to caption Refer to caption Refer to caption
$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},16^{2})$
Figure 3: Comparison between a theoretical anti-aliasing filter and Thera’s anti-aliasing, which is computed without convolutions or over-sampling. Note that the field was only supervised at $\Phi(\mathbf{x},1)$.
Figure 3 shows an example where we fit the signal at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 6. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, $\Phi(\mathbf{x},t)$ is known for any $t\geq t_{0}$.
By definition, such field $\Phi$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies. An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.
3.2 Thermal activations for super-resolution
The multi-scale signal representation defined above is a natural match for super-resolution, and particularly well-suited if arbitrary scale factors shall be covered. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture. Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed. To narrow down the (infinite) solution space to a unique result a prior must either be defined in an unsupervised fashion or learned from data. We describe below how these issues can be addressed, so that Thera can be used to model local fields that represent each pixel in the low-resolution image, a strategy initially proposed by LIIF [11] and used by most recent ASSR methods [22, 8, 41, 10].
Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size. Given a low-resolution image, the hypernetwork maps it to parameters for local fields, each spanning the area of one pixel of the low-resolution input. Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 41, 8, 10], followed by one or more layers which produce feature maps of depth $N$. In this work, we use a series of ConvNeXt [25] blocks for this purpose. The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and temporal coordinate $t$. The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values. Similar to LIIF, we use a local coordinate system for every field that is bounded in $[-0.5,0.5]\times[-0.5,0.5]$. The components $\textbf{W}_{1}$ are optimized directly rather than being an output of $\Psi$. Not only does this better fit the idea of having a consistent basis with which we represent each signal, but it also reduces the number of total parameters and, as we show later, leads to better results.
Figure 4: Overview of Thera. A hypernetwork estimates the parameter of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
The described scheme, Thera, is depicted in Fig. 4. It allows for arbitrary-scale super-resolution combining the power of proven feature extractors for SR and image restoration with the presented multiscale representation capabilities. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors from data that lie within the resolution range observed during training. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$. By training using multiple resolutions, we can also make $\kappa$ a trainable parameter to allow the network to adapt to different downsampling operators. We also manually set the bias of each channel of $\Phi$ to the RGB values of its associated low-resolution pixel.
Priors at $\bm{t=0}$. To perform ASSR for out-of-domain scaling factors, we can use unsupervised priors at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation, that prior takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well. We use an $\ell^{1}$ variant of TV:
$\mathcal{L}_{\text{TV}}(\Phi(\mathbf{x},0))=\mathbb{E}_{\mathbf{x}}[|\nabla\Phi(\mathbf{x},0)|].$ (7)
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than using a neighborhood approximation as done most often in previous work [36].
Implementation and training. We use JAX [7] to implement Thera to maximize the efficiency of the hypernetwork operations. A PyTorch [34] implementation to fit a Thera field to a single image is also provided. Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to a patch of size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with the Adam optimizer [21] with an initial learning rate of $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$ for $5\times 10^{6}$ iterations with a batch size of 16. Throughout the optimization, the learning rate is decayed to zero according to a cosine annealing schedule [26]. We use a mean absolute error (MAE) loss for training, to which the TV loss from Eq. 7 is added with a weight of $10^{-4}$. Standard data augmentations are used during training, including random flipping, rotations, and resizing of images. Analogous to previous work [40, 24, 41], we implement geometric self-ensembling, where the results of four rotated versions of the input are averaged at test time. We did not observe an improvement in performance if flips were included. Our code and pre-trained checkpoints will be made publicly available.
4 Results
Backbone (params.) Method Num. of In-distribution Out-of-distribution
add. params. $\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
EDSR- baseline [24] (1.22 M) MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
LIIF [11] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
LTE [22] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
CUF [41] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
CiaoSR [8] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
CLIT [10] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
Thera-S .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
Thera-M 0.48 M 34.86 31.21 29.22 26.94 23.87 22.32 21.31 20.60
Thera-L 1.41 M 34.88 31.22 29.24 26.96 23.89 22.33 21.32 20.61
RDN [48] (22.0 M) MetaSR [17] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
LIIF [11] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
LTE [22] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
CUF [41] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
CiaoSR [8] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
CLIT [10] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
Thera-S .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
Thera-M 0.48 M 34.97 31.39 29.43 27.15 24.05 22.49 21.44 20.71
Thera-L 1.41 M 34.98 31.40 29.45 27.16 24.06 22.49 21.45 20.72
SwinIR [23] (12.0 M) MetaSR [17] 0.45 M 35.15 31.40 29.33 26.94 23.80 22.26 21.26 20.54
LIIF [11] 0.35 M 35.17 31.46 29.46 27.15 24.02 22.43 21.40 20.67
LTE [22] 0.49 M 35.24 31.50 29.51 27.20 24.09 22.50 21.47 20.73
CUF [41] 0.30 M 35.31 31.56 29.56 27.23 24.10 22.52
CiaoSR [8] 1.43 M 35.28 31.53 29.56 27.26 24.12 22.52 21.48 20.59
CLIT [10] 15.7 M 35.29 31.55 29.55 27.26 24.11 22.51 21.45 20.70
Thera-S .008 M 35.27 31.57 29.58 27.25 24.13 22.55 21.52 20.77
Thera-M 0.48 M 35.29 31.58 29.59 27.27 24.15 22.56 21.52 20.77
Thera-L 1.41 M 35.29 31.59 29.61 27.30 24.17 22.59 21.54 20.79
Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the DIV2K validation set. The highest PSNR value per backbone and scaling factor is in bold and the second highest is underlined.
We evaluate the performance of Thera on a variety of datasets and with three popular feature extractors as backbones. We also conduct an ablation study that supports the design choices described in Sec. 3. Throughout this section, we will evaluate three variants of our method which differ solely in size of the hypernetwork and field:
  • Thera-S: We set the number of shared components to 32, and omit all but one layer of the hypernetwork, the layer that maps features to field parameters. This version adds only 8,256 parameters on top of the backbone.
  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork (5 ConvNeXt layers), approximately matching the parameter count of medium-sized competing methods (MetaSR [17], LIIF [11], LTE [22] and CUF [41]). This amounts to 0.48 M additional parameters.
  • Thera-L: We use 512 components and a larger hypernetwork (16 ConvNeXt layers), amounting to 1.41 M additional parameters. This version is comparable to CiaoSR [8] in terms of parameter count.
Datasets and metrics. Like previous methods, our models are trained with the DIV2K [1] training set, which is composed of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [47], BSDS100 [27] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [28] (high-resolution Japanese manga illustrations). Following prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Not all numbers could be computed for competing methods due to the lack of availability of code or checkpoints, see Appendix E. Additional quantitative metrics are also provided in the supplementary material.
Backbones. Following related work, we train our method jointly with three standard backbone feature extractors for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [48] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
4.1 Super-resolution performance
We begin with evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above. Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors. Thera achieves state-of-the-art performance. Thera-L outperforms all previous methods on 22 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline). Thera-M, which requires only a third as many additional parameters, still outperforms or – less frequently – is on par with competitors that have comparable parameter count in most cases. Notably, except for ×$\times 2$, Thera-M also matches the performance of more expensive, attention-based methods CiaoSR and CLIT, despite a more than 30×30$\times$30 × fewer parameters than CLIT.
Notably, our minimal version Thera-S – enabling arbitrary-scale SR with theoretical guarantees using only 8,192 additional parameters – regularly outperforms all comparison methods except for the most parameter-intensive ones when using the EDSR backbone. Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods (including attention-based ones) on 6 of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (and 170×170$\times$170 × fewer than CiaoSR) on top of the feature extractor. Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors. This confirms that having formal modeling of multi-scale representation baked into our network is highly beneficial, especially for higher scaling factors.
Method Set5 Set14 B100 Urban100 Manga109
$\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4
MetaSR 38.26 34.77 32.47 34.14 30.66 28.85 32.39 29.31 27.75 33.29 29.12 26.76 39.46 34.62 31.37
LIIF 38.28 34.87 32.73 34.14 30.75 28.98 32.39 29.34 27.84 33.36 29.33 27.15 39.57 34.68 31.71
LTE 38.33 34.89 32.81 34.25 30.80 29.06 32.44 29.39 27.86 33.50 29.41 27.24 39.63 34.79 31.79
CUF 38.38 34.92 32.83 34.33 30.84 29.05 32.47 29.40 27.88 33.65 29.55 27.32
CiaoSR 38.38 34.91 32.84 34.33 30.82 29.08 32.47 29.42 27.90 33.65 29.52 27.42 39.67 34.84 31.91
CLIT 38.41 34.97 32.86 34.27 30.85 29.08 32.46 29.42 27.91 33.56 29.43 27.25
Thera-S 38.32 34.88 32.75 34.30 30.85 29.07 32.46 29.39 27.88 33.61 29.56 27.25 39.62 34.93 31.91
Thera-M 38.36 34.90 32.80 34.30 30.86 29.12 32.47 29.40 27.89 33.58 29.56 27.27 39.70 35.02 32.04
Thera-L 38.41 34.93 32.82 34.33 30.87 29.11 32.49 29.42 27.91 33.67 29.62 27.33 39.71 35.01 32.07
Table 2: Evaluation on common benchmark datasets for in-distribution scale factors, with a SwinIR backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
We further report the performance of our method on the five considered benchmark datasets in Tab. 2. For these experiments, we use the SwinIR backbone. Overall, due to the out-of-distribution nature of the data and the small size of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent. This suggests that future work on ASSR would benefit from using larger-scale datasets in the future. Nonetheless, our methods perform best in 12 out of 15 cases. We once again observe that the performance of Thera-S is often comparable to that of methods that require orders of magnitude more parameters.
In our qualitative analysis – an excerpt of which is given in Fig. 5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, in particular with repeating structure. This once again highlights the advantages of using an accurate observation model in our field formulation.
DIV2K
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Set14
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Manga109
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution. We use the SwinIR [23] backbone for all methods. More examples are given in Appendix B.
4.2 Multi-scale consistency
We are interested in learning a super-resolution model that is consistent across scales. For instance, if we super-resolve an image with a scaling factor of $\times$4 then downsample it by a factor of $\times$2, we expect the result to be very similar to what we get by simply super-resolving the original image by a factor of $\times$2. We report in Tab. 3 the quality of this match for several scaling factor pairs for Thera – which uses a formal observation model – and for three comparison methods that learn such model from data. First, we notice that better super-resolution quality does not correlate with scale consistency – e.g., LTE achieves better consistency than CiaoSR for most scaling factor pairs. Second, while other methods generally exhibit good consistency, even for out of distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
$(2,4)$ $(2,6)$ $(2,12)$ $(2,24)$ $(3,12)$ $(3,24)$ $(4,24)$
LIIF [11] 44.88 43.75 42.21 41.48 43.64 42.52 44.20
LTE [22] 44.10 44.40 44.37 44.40 48.33 48.48 51.42
CiaoSR [8] 44.24 43.31 42.10 41.56 41.47 41.03 45.47
Thera-S 45.79 46.67 46.80 46.89 52.38 52.66 54.84
Thera-M 45.26 46.10 46.34 46.41 51.91 52.16 54.64
Thera-L 45.42 46.24 46.43 46.47 51.56 51.77 54.33
Table 3: Comparison of multi-scale consistency (in PSNR) for various combinations of scaling factors.
4.3 Ablation studies
In Tab. 4, we ablate individual components and design choices of our method to understand their contribution to the overall performance. These comparisons are done using Thera-M using the EDSR-baseline backbone.
Single scale training. We ran three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization. $\kappa$ was fixed for these experiments since it requires multiscale training to be optimized. We observe equal or even superior performance compared to our base case for the scale used for training (marked in yellow in Tab. 4), but we see a significant drop in generalization.
Trainable $\kappa$. We observe a small drop in performance when using the fixed, theoretically derived value for $\kappa$ (Eq. 5). This suggests that there are still effects that have not been accounted for in our model, but the effects are very minor.
Geometric self-ensemble. As observed by previous work [40, 24, 41], we see a notable boost in performance when using geometric self-ensemble (GSE). Nevertheless, if a user prioritizes speed over quality, this addition can be disabled without retraining the network at test-time.
Total variation. We observe negligible effects for in-domain scaling factors when no TV is used, but performance degrades significantly for out-of-distribution scales.
No thermal activations. We replace thermal activations (Eq. 4) for standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A loss in performance shows the value of our proposed thermal activation for multiscale representation.
Shared components. We observe a noticeable performance drop in all scales if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$. Despite being even more flexible, it seems challenging to learn per-field components and other correlated parameters simultaneously.
Method In-distribution Out-of-distribution
$\times$2 $\times$3 $\times$4 $\times$6 $\times$18 $\times$30
Thera-M 34.86 31.21 29.22 26.94 22.33 20.60
$\times$2 only, fixed $\kappa$ 34.98 30.96 28.81 26.40 21.93 20.30
$\times$3 only, fixed $\kappa$ 34.71 31.22 29.22 26.88 22.24 20.53
$\times$4 only, fixed $\kappa$ 34.44 31.14 29.22 26.95 22.30 20.58
Fixed $\kappa$ 34.85 31.20 29.22 26.93 22.30 20.58
No GSE 34.79 31.14 29.16 26.88 22.28 20.56
No TV prior 34.87 31.21 29.21 26.86 20.82 19.12
No $\xi$ 34.80 31.12 29.13 26.87 22.28 20.58
Non-shared comps. 34.80 31.13 29.14 26.87 22.28 20.57
Table 4: Ablations have been run using Thera-M using the EDSR-baseline backbone, a representative case for all other settings.
4.4 Limitations and future work
Thermal activations as introduced in this paper, and by extension our proposed ASSR method, come with relatively strict architectural requirements that only allow for a single hidden layer in the neural field. While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder. While we have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work. This could result in higher parameter efficiency, and potentially better generalization. There may also be other formulations with similar characteristics which do not require the usage of a hypernetwork, which may allow for more streamlined training.
We noticed a relative drop in performance w.r.t. other methods for $\times$2 SR. We speculate that our weakness relative to higher scaling factors is due to this scale being close to the interface between continuous (field) and grid (pixel) representations, which Thera’s formalism does not model. Further exploration on this transition regime for such hybrid structures would be useful to better understand and use these hybrid representations. This could also benefit Instant-NGP [31] and related methods.
Finally, we have only explored TV as a method for regularizing $\Phi$ at $t=0$, but we expect other signal priors to be even more effective. Priors at $t=0$, possible only by using Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.
5 Conclusion
We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations. Our method, Thera, implicitly describes an image as a combination of sinusoidal components which can be modulated based on their frequency to perform Gaussian filtering (anti-aliasing) analytically with little overhead. Our experimental evaluation shows that Thera is superior to competing methods in most cases, while also being more parameter efficient and providing theoretical guarantees. We also hope to inspire others to apply Thera to other problems where it is likely to be beneficial due to its Gaussian observation model, such as demosaicing, burst super-resolution, and neural radiance fields.
Supplementary Material
Appendix A Theory
A.1 Preliminaries
As was described in Sec. 3.1, the idea is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
$\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;.$ (8)
The reason for this is that the analytical solution to the (isotropic) heat equation can be modelled as a convolution of the initial state $\Phi(\mathbf{x},0)$ with a Gaussian kernel
$g(\mathbf{x},t)=\frac{1}{4\pi\kappa t}\cdot\exp\left(-\frac{x_{1}^{2}+x_{2}^{2}}{4\kappa t}\right).$ (9)
By fitting the data at $\Phi(\mathbf{x},1)$, we are assuming a Gaussian point spread function (PSF) with the shape
$\text{PSF}(\mathbf{x})=\frac{1}{4\pi\kappa}\cdot\exp\left(-\frac{x_{1}^{2}+x_{2}^{2}}{4\kappa}\right).$ (10)
In this formulation, we attempt to recover a “pure” signal at $t=0$ or higher sampling rates $10<t<$ 1 given an observation at $t=1$. Note that
$\Phi(\mathbf{x},t)\text{ is}\begin{cases}\text{meaningless},&\text{if}\ t<0\\ \text{pure signal},&\text{if}\ t=0\\ \text{ill-posed},&\text{if}\ 0<t<1\\ I,&\text{if}\ t=1\\ \text{well-posed},&\text{if}\ t>1\\ \end{cases}$ (11)
The ill-posed problem for $10<t<$ 1 is the interesting case. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
For all the formulations here, we define the image $I$ to correspond to the coordinates $x_{1},x_{2}\in[-0.5,0.5]$.
A.2 Thermal diffusivity coefficient
To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$. One way to do so is to match the cut-off frequency of the filter in Eq. 9 at $t=1$ to the well-known Nyquist frequency given by the sampling rate. We take the cut-off frequency of the Gaussian filter defined in Eq. 9 to be the frequency whose amplitude is halved, which is
$f_{c}=\sqrt{\ln(4)}\cdot\sigma_{f}=\frac{\sqrt{\ln(4)}}{2\pi\sqrt{2kt}}$ (12)
For the signal compressed into the domain $[-0.5,0.5]$, we can compute the Nyquist frequency to be
$f_{\text{Nyquist}}=\frac{N}{2},$ (13)
where $N$ is the number of samples along a given dimension. This formulation assumes even sampling over $x_{1}$ and $x_{2}$.
If we solve for $f_{c}=f_{\text{Nyquist}}$ at $t=1$, we get
$\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}N^{2}}.$ (14)
For the Thera formulation presented in Sec. 3.2, we want $\Phi$ to contain a single pixel at $t=1$, therefore we initialize $\kappa$ with
$\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}}.$ (15)
Note that the exact value of $\kappa$ will depend on the exact characteristics of the system that is being modelled. Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but is also more prone to aliasing.
Finally, we would like to highlight that Eq. 12 is specific to the case where $\mathbf{x}$ is 2-dimensional. The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using fields in spaces with numbers of spatial dimensions other than 2. Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the desired formulas for a Gaussian filter with the appropriate number of dimensions.
A.3 Relationship between $t$ and $s$
Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate. If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it with $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$. Using Eq. 12 and Eq. 13, we can easily find the quadratic relationship
$t=S^{2}.$ (16)
For instance, if we want to upsample the image by a factor of 2, we should use
$t=0.5^{2}=0.25.$ (17)
Thus, $10<t<$ 1 refer to super-resolution, while $t>1$ refers to downsampling.
In-distribution Out-of-distribution
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$$\times 12$$\times 18$$\times 24$$\times 30$
LIIF [11] .942 .883 .830 .753 .702 .641 .594 .568 .551
LTE [22] .942 .883 .832 .755 .704 .643 .595 .569 .552
CiaoSR [8] .942 .883 .832 .756 .706 .645 .598 .572 .551
Thera-S .943 .884 .833 .757 .706 .644 .596 .569 .552
Thera-M .943 .885 .834 .758 .707 .645 .596 .569 .552
Thera-L .942 .885 .834 .759 .708 .646 .597 .570 .552
Table 5: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the SwinIR [23] backbone for all models.
Method Set5 Set14 B100 Urban100 Manga109
$\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12
MetaSR 29.09 27.02 24.82 26.58 25.09 23.33 25.94 24.87 23.59 24.16 22.75 21.31 27.29 24.96 22.35
LIIF 29.46 27.36 26.82 25.34 26.07 25.01 24.59 23.14 27.69 25.28
LTE 29.50 27.35 26.86 25.42 26.09 25.03 24.62 23.17 27.83 25.42
CUF 29.57 26.91 26.11 24.69
CiaoSR 29.62 27.45 24.96 26.88 25.42 23.38 26.13 25.07 23.68 24.84 23.34 21.60 28.01 25.61 22.79
CLIT 29.69 27.62 26.94 25.55 26.15 25.09 24.77 23.33
Thera-S (ours) 29.45 27.33 25.00 26.85 25.40 23.40 26.11 25.05 23.69 24.68 23.23 21.55 27.94 25.49 22.73
Thera-M (ours) 29.47 27.43 25.07 26.90 25.41 23.40 26.13 25.07 23.69 24.70 23.25 21.57 28.05 25.59 22.79
Thera-L (ours) 29.48 27.46 25.09 26.94 25.45 23.42 26.14 25.08 23.71 24.79 23.32 21.62 28.13 25.69 22.85
Table 6: Evaluation on common benchmark datasets for out-of-distribution scale factors, using the SwinIR backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
A.4 Other filters
In theory, the formulation presented in Sec. 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely. Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation. Initial explorations of a sharp low-pass filter which completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely to how this also affected gradients during training. It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way. For quantitative evaluations this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios this may be desirable.
A.5 Initialization of components
We have noticed during our experiments that the initialization of the components ($\mathbf{W}_{1}$ in Eq. 2) is very important. Sitzmann et al. [38] made similar observations when periodic activation functions were first used for neural fields. The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training. Thus, we chose to initialize $\mathbf{W}_{1}$ such that
$p(\left|\nu(\mathbf{w}_{1})\right|)\propto\left|\nu(\mathbf{w}_{1})\right|$ (18)
up to a given maximum frequency, allotting more components to higher frequencies. See code for more details.
Appendix B Additional qualitative results
Figure 8 and Fig. 9 contain additional qualitative results for lower and higher scaling factors, respectively. Again, we see that Thera can reconstruct the targets more faithfully. This is particularly visible in repetitive structural elements. We suspect that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this.
Appendix C Additional quantitative results
Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab. 1. We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric. Furthermore, Tab. 6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$). Finally, in Fig. 7 we show the number of additional parameters for various methods on the ×$\times 3$ upsampling task.
Appendix D Hypernetwork architecture
Figure 6: Schematics of the hypernetworks for a) Thera-S, b) Thera-M and c) Thera-L.
In Fig. 6, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L. The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image. It is then followed by a number of (residual) ConvNeXt [25] blocks. Each block consists of i) a $3\times 3$ convolution that maps to $d$-dimensional features, ii) a layer normalization [2] operation, iii) a $1\times 1$ convolution with $4d$ filters, iv) a GELU [15] non-linearity, and v) another $1\times 1$ convolution with $d$ filters. Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al. [25] – to $3\times 3$ to reduce the number of parameters. Furthermore, note that in Thera-L, the first convolution per block uses $d$ groups (effectively implementing a channel-wise convolution), while for the two smaller variants we used full convolutions, i.e., a single group. In our implementation, Thera-S uses zero ConvNeXt blocks, Thera-M uses 5 blocks with $d=64$, and Thera-L uses 6 blocks with $d=64$ followed by 7 blocks with $d=96$ and 3 blocks with $d=128$. For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [38]. There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L. Let $c$ be the number of components used (32, 512 and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork. The total parameter count of the hypernetwork is 8,192, 0.48 M and 1.41 M for the three variants, respectively. To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters. We highlight that these do not come from the hypernetwork.
Appendix E Replication of other methods
We encountered challenges attempting to recreate results of some of the competitor methods, which explains why some numbers are missing in various tables or differ from the originally reported ones. Details are provided below.
CUF [41]. At the time of writing, there was no public code or checkpoints available for CUF. Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper. We have denoted those missing values with “—” in the tables. Furthermore, we could not create any qualitative samples and could not include CUF in the multi-scale consistency study in Sec. 4.2.
CLIT [10]. For CLIT, at the time of writing, there is public code available, however, checkpoints are not. We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU hours: over 1 month using 8×8$\times$8 × Nvidia GeForce RTX 3090s. The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
CiaoSR [8]. We noticed that CiaoSR’s border cropping prior to evaluation on the DIV2K dataset in the official implementation deviated slightly from the rest of the comparison methods. We aligned this to align the evaluation of all methods and re-computed all DIV2K numbers. This explains the slight deviation from the original paper.
Figure 7: Plots depicting the number of additional parameters vs. PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 3$ upsampling task.
DIV2K ×$\times 2$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DIV2K ×$\times 3$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DIV2K ×$\times 3$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100 ×$\times 2$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100 ×$\times 3$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Manga109 ×$\times 3$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Manga109 ×$\times 3$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
Figure 8: Further qualitative examples for ×$\times 2$ and ×$\times 3$ upsampling factors. Best viewed zoomed in.
DIV2K ×$\times 4$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DIV2K ×$\times 4$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100 ×$\times 4$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100 ×$\times 4$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Urban100 ×$\times 6$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Manga109 ×$\times 6$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Manga109 ×$\times 6$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
Figure 9: Further qualitative examples for ×$\times 4$ and ×$\times 6$ upsampling factors. Best viewed zoomed in.

### v2

@@
 Abstract
-Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at different rates.
-However, in such formulation, the point-wise query of field values does not naturally match the point spread function (PSF) of a given pixel.
-In this work we present a novel way to design neural fields such that points can be queried with a Gaussian PSF, which serves as anti-aliasing when moving across resolutions for ASSR.
-We achieve this using a novel activation function derived from Fourier theory and the heat equation.
-This comes at no additional cost: querying a point with a Gaussian PSF in our framework does not affect computational cost, unlike filtering in the image domain.
-Coupled with a hypernetwork, our method not only provides theoretically guaranteed anti-aliasing, but also sets a new bar for ASSR while also being more parameter-efficient than previous methods.
+Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at arbitrary rates.
+However, the point-wise query of the neural field does not naturally match the point spread function (PSF) of a given pixel, which may cause aliasing in the super-resolved image.
+We present a novel way to design neural fields such that points can be queried with an adaptive Gaussian PSF, so as to guarantee correct anti-aliasing at any desired output resolution.
+We achieve this with a novel activation function derived from Fourier theory.
+Querying points with a Gaussian PSF, compliant with sampling theory, does not incur any additional computational cost in our framework, unlike filtering in the image domain.
+With its theoretically guaranteed anti-aliasing, our method sets a new state of the art for ASSR, while being more parameter-efficient than previous methods.
+Notably, even a minimal version of our model still outperforms previous methods in most cases, while adding 2-4 orders of magnitude fewer parameters.
+Code and pretrained models are available at https://github.com/prs-eth/thera.
 1 Introduction
 For years, learning-based image super-resolution (SR) methods have achieved ever better results.
-But unlike interpolation methods, which allow us to resample a given image at any resolution, such methods require retraining for each desired scaling factor.
-In recent years, several supervised methods for arbitrary-scale SR (ASSR) have been proposed to allow end-users to choose any scaling factor, rendering such systems much more flexible [17].
-Notably, Chen et al.
-[11] pioneered the use of neural fields for single-image super-resolution (SR) with LIIF, exploiting the continuous representation to enable SR at arbitrary scaling factors.
+But unlike interpolation methods, which allow us to resample a given image at any resolution, those methods require retraining for each desired scaling factor.
+Recently, several schemes for arbitrary-scale SR (ASSR) have been proposed that allow end-users to choose any scaling factor, rendering the method much more flexible [17].
+Notably, with LIIF, Chen et al.
+[11] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors.
 Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10].
 This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [46].
-By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and analytically differentiable.
+By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
 Among others, they have been widely adopted for 3D reconstruction with radiance [32], occupancy [31, 37], and distance fields [35, 39].
 Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components.
-We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies.
+We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies (insets illustrate components decaying at different rates).
 This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
 Neural fields offer a compelling representation of continuous signals.
-However, for many tasks it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3].
-This is not trivial [3, 4, 16, 5].
-For ASSR, where the signal should be observed at various scales and is prone to aliasing, a neighborhood-dependent observation model is desirable.
-Using appropriate observation modeling has not only the potential to reduce aliasing but also to more accurately reflect the physical processes of image acquisition.
+However, for many tasks, it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3].
+ASSR calls for a neighborhood-dependent observation model, as the signal should be observed at various sampling frequencies and is prone to aliasing.
+Incorporating a physically plausible observation model is not trivial [3, 4, 25, 49, 16, 5], but has the potential to avoid aliasing.
 For this reason, Chen et al.
-[11] already took a first step towards learning multiscale representations through cell encoding.
-For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the location $\mathbf{x}$ as a tool to control different frequencies in $\Phi$.
+[11] already took a first step towards learning multiscale representations, via cell encoding.
+For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the input coordinates, as a means to control different frequencies in the field.
+Lindell et al.
+[25] and Yang et al.
+[49] have succeeded in performing subband decomposition of neural fields, though bands are architecturally fixed to a small number of discrete intervals and cannot be modified after training.
 In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction.
-Furthermore, this filtering incurs no additional overhead: The cost of querying a field’s value does not depend on the size of the anti-aliasing filter.
-The key insight is that when using sinusoidal activation functions [38], one can decay different components at different rates based on their frequencies, according to Fourier theory.
-Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, an ASSR method with neural fields that uses a novel thermal activation function (see Fig.
-1).
-This formulation makes it possible to analytically compute Gaussian-blurred versions of a field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius.
-This is equivalent to querying $\Phi$ with a Gaussian PSF.
-Thera outperforms competing methods in most of our experiments, is far more parameter-efficient, and provides theoretical guarantees w.r.t.
-anti-aliasing and multi-scale representation.
-We demonstrate that a lightweight variant of our algorithm still outperforms most previous methods while using 2-4 orders of magnitude fewer parameters.
-To provide theoretical guarantees for our thermal activations, certain constraints are imposed on the field’s neural architecture.
-Fortunately, these constraints can easily be overcome with the help of a hypernetwork [14], so that our method can seamlessly integrate a wide range of well-proven SR backbones.
-In summary, our main contributions are:
+Interestingly, this filtering incurs no additional overhead: The cost of querying field values is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
+The key insight is that sinusoidal activation functions [40] offer the possibility to selectively decay each component according to its spatial frequency, following Fourier theory.
+Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, a neural field method for ASSR underpinned by a novel thermal activation function, see Fig.
+1.
+That formulation makes it possible to analytically compute Gaussian-blurred versions of the neural field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius.
+I.e., one can query $\Phi$ with a Gaussian PSF that is matched to the output resolution to prevent aliasing.
+[^1] While our framework is in principle not bound to the heat equation or Gaussian PSFs, it is an intuitive interpretation for readers familiar with signal theory.
+Thera provides theoretical guarantees w.r.t.
+anti-aliasing and multi-scale representation, in exchange for certain constraints on the neural architecture that can easily be overcome with a hypernetwork [14].
+Empirically, it outperforms competing methods in most of our experiments and is far more parameter-efficient.
+We demonstrate that a lightweight variant of our model still outperforms most previous methods while adding 2-4 orders of magnitude fewer parameters.
+To the best of our knowledge, Thera is the first neural field method to allow bandwidth control at test time.
+Our main contributions are:
   1. 1.
-    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and as a consequence anti-aliasing with minimal overhead.
+    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and therefore anti-aliasing with minimal overhead.
   2. 2.
-    We use thermal activations to build Thera, a novel method for ASSR which is superior to previous methods while being more parameter efficient and providing theoretical guarantees w.r.t.
-multiscale representation.
+    We use thermal activations to build Thera, a novel method for ASSR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
 2 Related Work
-Arbitrary-scale SR.
-In recent years, research on arbitrary-scale super-resolution (ASSR) methods has gained attention.
+2.0.1 Arbitrary-scale SR.
 In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods.
-The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork, conditioned on the desired scaling factor.
-An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by a neural field that is locally conditioned with features extracted from the low-resolution image.
+The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork [14] conditioned on the desired scaling factor.
+An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by local neural fields that are conditioned with features extracted from the low-resolution input image.
 The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution.
 Following work has since built upon the LIIF framework.
-For instance, UltraSR [45] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields.
+For instance, UltraSR [48] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields [32, 3, 4, 5].
 LTE [22] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP.
@@
 These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data.
-In this paper, we explore a way to directly integrate such a model into the neural field representation.
+In this paper, we explore a way to directly integrate a (Gaussian) observation model into the neural field representation.
 Implicit neural representations.
-A neural field is a neural network trained to map real-valued coordinates onto a field’s values.
-Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 38, 39, 11, 22, 12], 3D scenes (e.g., represented as signed distance fields [33, 38, 37, 42], occupancy fields [29, 35], LiDAR fields [19], view-dependent radiance fields [30, 3, 4, 5]), or digital humans [46, 49, 9, 44, 13].
+A neural field is a neural network trained to map coordinates onto values of some physical quantity.
+Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 40, 41, 11, 22, 12, 45], 3D scenes (e.g., represented as signed distance fields [35, 40, 39, 44, 45], occupancy fields [31, 37], LiDAR fields [19], view-dependent radiance fields [32, 3, 4, 5, 45]), or digital humans [50, 53, 9, 47, 13].
 We refer the interested reader to Xie et al.
@@
 Frequently, it is desirable to impose some prior over the space of learnable representations.
-A common approach for such conditioning of neural fields is encoder-based inference [43], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 41, 8, 10].
-The encoded latent variables $\bm{z}$ are then used to condition field parameters $\bm{\theta}$, for instance by simply concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork  [14].
-The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, mapping latent codes to field parameters.
-An early example of this is Sitzmann et al.
-[38], and this approach is gaining popularity [43].
+A common approach for such conditioning is encoder-based inference [46], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 43, 8, 10].
+The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [14].
+The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, directly mapping latent codes to parameters of the field $\bm{\theta}$.
+An early example of this approach, which is gaining popularity [46], was proposed by Sitzmann et al.
+[40].
 Anti-aliasing in neural fields.
@@
 [5] tackle anti-aliasing for the increasingly popular Instant-NGP [33] approach.
+Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [25] or polynomial neural fields [49], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks.
+Such representations would not be a good fit for ASSR since they do not allow for continuous anti-aliasing, nor bandwidth control at test time.
 In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
 3 Method
-To understand how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis.
-While we stick to the 2D case here for simplicity, all theory is valid for signals with any number of dimensions.
-We refer the reader to [32] for an in-depth analysis of sampling and aliasing.
-To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ in 2D image space, $\mathbf{x}\in\mathbb{R}^{2}$, it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate.
-One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, and then sample the band-limited signal $f\circledast g(\mathbf{x})$.
-Here, we show that one can naturally integrate this behavior into a neural field’s architecture.
-The key to this is knowing that if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor.
+To expound how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis.
+For simplicity, we stick to 2-dimensional signals $\mathbf{x}\in\mathbb{R}^{2}$, but all theory is valid for any number of dimensions.
+We refer the reader to [34] for a comprehensive discussion of sampling and aliasing.
+To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate.
+One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$.
+We will show that one can naturally integrate this behavior into a neural field’s architecture.
+The key to this is that, if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor.
 This equivalence is depicted in Fig.
 2.
-While decomposing a given signal into sinusoidal components is normally a costly operation, we show that we can avoid it by designing a neural field in which the decomposition itself is the native representation of the signal.
+Decomposing a given signal into sinusoidal components is normally a costly operation, but it turns out that we can avoid it by designing the neural field such that the decomposition itself is the native representation of the signal.
 3.1 Thermal Activation Functions
 The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation.
-Other implicit representation methods, such as SIREN [38], optimize the parameters $\bm{\theta}_{\Phi}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$.
-Our key insight is the following: What if we take this field to represent the temperature distribution on a surface at a given time $t_{0}$?
-It is well known how to solve the heat equation forwards in time – its solution (on an isotropic surface) is given by a Gaussian blur of the state at $t_{0}$, where the blur radius depends on the time difference $t-t_{0}$.
+Other implicit representation methods, such as SIREN [40], optimize the parameters $\bm{\theta}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$.
+Our key insight is that we can interpret this field as the temperature distribution on an (isotropic) surface at a given time $t_{0}$.
+It is well known how to solve the heat equation forwards in time – its solution is obtained by blurring the initial state at $t_{0}$ with a Gaussian kernel whose radius depends on the time difference $t-t_{0}$.
 Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
 $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;,$ (1)
-where $\kappa>0$ is the thermal diffusivity of the medium.
-The solution for $0\leq t<t_{0}$ is, however, ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
+with $\kappa>0$ the thermal diffusivity of the medium.
+In contrast, the solution for $0\leq t<t_{0}$ is ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
 Figure 2: If we have access to the sinusoidal decomposition of a signal, filtering can be done analytically by scaling each component appropriately.
 We design our neural field so that we have access to such a decomposition at no extra cost.
-A naive attempt to ensure that $\Phi$ respects the heat equation in Eq.
-1 would be to include an additional loss term: $\left|\frac{\partial\Phi}{\partial t}-\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi\right|$.
-Despite our best efforts, this approach was unsuccessful.
-Nevertheless, it turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq.
-1.
-We do so by formulating $\Phi$ as a two-layer perceptron
-$\mathbf{y}=\mathbf{W}_{2}\cdot\xi(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t)+\textbf{b}_{2},$ (2)
-where $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$, described in more detail later; and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time.
-We set $t=0$ to be the (unobservable) initial thermal state of a given system represented as a sum of sinusoidal components:
+It turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
+$\Phi(\mathbf{x},t)$ $=\mathbf{W}_{2}\cdot\xi\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t\right)+\textbf{b}_{2},$ (2)
+where $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$ are MLP parameters, $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$ (described in more detail later), and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time.
+We set $t=0$ to be the initial thermal state of a given system represented as a sum of sinusoidal components:
 $\xi(\mathbf{z},\nu,\kappa,0)=\sin(\mathbf{z}).$ (3)
@@
 $\xi(\mathbf{z},\nu,\kappa,t)=\sin(\mathbf{z})\cdot\exp(-\left|\nu\right|^{2}\kappa t).$ (4)
-To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally over time, regardless of their orientation).
+To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally, regardless of their orientation).
 Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
@@
 $t=S^{2}.$ (6)
-The derivations of these values can be found in Appendix A.
+The derivations of these values can be found in Appendix 0.A.
 In our formulation, $\mathbf{W}_{1}$ can be seen as a frequency bank whose components (“basis functions”) make up the signal $\Phi(\mathbf{x},0)$, and the vector $\mathbf{b}_{1}$ represents the phase shifts of those components.
@@
 Image domain
-Refer to caption Refer to caption Refer to caption
-Original Anti-aliasing for $\times$1/4 Anti-aliasing for $\times$1/16
-Thera
-Refer to caption Refer to caption Refer to caption
-$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},16^{2})$
-Figure 3: Comparison between a theoretical anti-aliasing filter and Thera’s anti-aliasing, which is computed without convolutions or over-sampling.
-Note that the field was only supervised at $\Phi(\mathbf{x},1)$.
-Figure 3 shows an example where we fit the signal at $t=1$ to the image.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA $\times$1/32 AA
+Neural field
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$ $\Phi(\mathbf{x},32^{2})$
+Figure 3: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with thermal activations (bottom row), which is computed without convolutions or over-sampling.
+The neural field was supervised only at $\Phi(\mathbf{x},1)$.
+Figure 3 shows an example where we fit the field at $t=1$ to the image.
 After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq.
 6.
-We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, $\Phi(\mathbf{x},t)$ is known for any $t\geq t_{0}$.
-By definition, such field $\Phi$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies.
+We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
+By definition, our field $\Phi(\mathbf{x},t)$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies.
+Note that modulating deeper fields, as is done in [30], would no longer guarantee anti-aliasing.
 An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.
@@
 Still, two challenges must be addressed.
-First, our formulation restricts the choice of architecture.
+First, our formulation restricts the choice of architecture to MLPs with a single hidden layer.
 Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed.
-To narrow down the (infinite) solution space to a unique result a prior must either be defined in an unsupervised fashion or learned from data.
-We describe below how these issues can be addressed, so that Thera can be used to model local fields that represent each pixel in the low-resolution image, a strategy initially proposed by LIIF [11] and used by most recent ASSR methods [22, 8, 41, 10].
+To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data.
 Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size.
-Given a low-resolution image, the hypernetwork maps it to parameters for local fields, each spanning the area of one pixel of the low-resolution input.
+The hypernetwork maps the low-resolution input image to parameters of local fields, each spanning the area of one pixel of the low-resolution input as originally proposed by LIIF [11] and adopted by most recent ASSR methods [22, 8, 43, 10].
 Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field.
-The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 41, 8, 10], followed by one or more layers which produce feature maps of depth $N$.
-In this work, we use a series of ConvNeXt [25] blocks for this purpose.
-The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and temporal coordinate $t$.
+The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 43, 8, 10], followed by one or more layers that produce feature maps of depth $N$.
+In this work, we use a series of ConvNeXt [26] blocks.
+The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$.
 The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values.
-Similar to LIIF, we use a local coordinate system for every field that is bounded in $[-0.5,0.5]\times[-0.5,0.5]$.
-The components $\textbf{W}_{1}$ are optimized directly rather than being an output of $\Psi$.
-Not only does this better fit the idea of having a consistent basis with which we represent each signal, but it also reduces the number of total parameters and, as we show later, leads to better results.
-Figure 4: Overview of Thera. A hypernetwork estimates the parameter of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
-The described scheme, Thera, is depicted in Fig.
+Similar to LIIF, the local coordinate system for every field is bounded by $[-0.5,0.5]\times[-0.5,0.5]$.
+In practice, we directly optimize a single global frequency bank $\textbf{W}_{1}$, rather than having the hypernetwork predict a separate $\textbf{W}_{1}$ for each low-resolution pixel.
+Not only does this better fit the idea of to represent the signal with a single, consistent basis; it also reduces the total parameter count and empirically leads to better results.
+Figure 4: Overview of Thera. A hypernetwork estimates the parameters of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
+The described scheme, which we call Thera, is depicted in Fig.
 4.
-It allows for arbitrary-scale super-resolution combining the power of proven feature extractors for SR and image restoration with the presented multiscale representation capabilities.
-As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors from data that lie within the resolution range observed during training.
+It allows for arbitrary-scale super-resolution, combining the proposed multi-scale signal representation with the expressivity of proven feature extraction backbones for SR and image restoration.
+As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data.
 E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$.
-By training using multiple resolutions, we can also make $\kappa$ a trainable parameter to allow the network to adapt to different downsampling operators.
-We also manually set the bias of each channel of $\Phi$ to the RGB values of its associated low-resolution pixel.
-Priors at $\bm{t=0}$.
-To perform ASSR for out-of-domain scaling factors, we can use unsupervised priors at $t=0$.
+By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators.
+Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel.
+Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.
+Prior at $\bm{t=0}$.
+To perform ASSR for out-of-domain scaling factors, we can use an unsupervised prior at $t=0$.
 Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods.
-In our implementation, that prior takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well.
+In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well.
 We use an $\ell^{1}$ variant of TV:
 $\mathcal{L}_{\text{TV}}(\Phi(\mathbf{x},0))=\mathbb{E}_{\mathbf{x}}[|\nabla\Phi(\mathbf{x},0)|].$ (7)
-Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than using a neighborhood approximation as done most often in previous work [36].
+Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than falling back to a neighborhood approximation as in most previous work [38].
 Implementation and training.
-We use JAX [7] to implement Thera to maximize the efficiency of the hypernetwork operations.
+We have implemented Thera in JAX [7] to maximize the efficiency of the hypernetwork operations.
 A PyTorch [36] implementation to fit a Thera field to a single image is also provided.
-Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to a patch of size $48^{2}$.
+Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$.
 As corresponding targets, $48^{2}$ random pixels are sampled from the target patch.
-We train with the Adam optimizer [21] with an initial learning rate of $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$ for $5\times 10^{6}$ iterations with a batch size of 16.
-Throughout the optimization, the learning rate is decayed to zero according to a cosine annealing schedule [26].
-We use a mean absolute error (MAE) loss for training, to which the TV loss from Eq.
+We train with the Adam optimizer [21] with batch size $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$.
+The learning rate is decayed to zero according to a cosine annealing schedule [27].
+As reconstruction loss, we minimize the mean absolute error (MAE), to which the TV loss from Eq.
 7 is added with a weight of $10^{-4}$.
-Standard data augmentations are used during training, including random flipping, rotations, and resizing of images.
-Analogous to previous work [40, 24, 41], we implement geometric self-ensembling, where the results of four rotated versions of the input are averaged at test time.
-We did not observe an improvement in performance if flips were included.
-Our code and pre-trained checkpoints will be made publicly available.
+Training data is augmented by standard data augmentations, namely random flipping, rotation, and resizing.
+As in previous work [42, 24, 43], we implement geometric self-ensembling, where the results for four rotated versions of the input are averaged at test time.
+We did not observe a performance gain when also including reflections.
 4 Results
+We evaluate Thera on a variety of datasets and with three popular feature extractors as backbones.
+We also conduct an ablation study that supports the design choices described in Sec.
+3.
+Throughout this section, we will evaluate three variants of our method which differ solely in the size of the hypernetwork and the number of field parameters:
+  • Thera-S: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and with only a single layer in hypernetwork, which maps features to field parameters.
+This version adds only 8,256 parameters on top of the backbone.
+  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork with 5 ConvNeXt layers.
+Its parameter count of 0.48 M approximately matches that of medium-sized competitors (MetaSR [17], LIIF [11], LTE [22] and CUF [43]).
+  • Thera-L: A large version with 512 components and a deeper hypernetwork of 16 ConvNeXt layers, for a total of 1.41 M additional parameters, comparable to CiaoSR [8].
+Datasets and metrics.
+Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes.
+We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [51], BSDS100 [28] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [29] (high-resolution Japanese manga illustrations).
+Like prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets.
+Additional metrics are given in Appendix 0.C.
+Not all numbers could be computed for competing methods, because some do not provide public code or checkpoints, see Appendix 0.E.
+Backbones.
+Following related work, we combine our method with three standard backbones for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [52] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
+Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the DIV2K validation set.
+The highest PSNR value per backbone and scaling factor is in bold and the second highest is underlined.
 Backbone (params.) Method Num. of In-distribution Out-of-distribution
@@
 Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
-EDSR- baseline [24] (1.22 M) MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
+EDSR- baseline [24] (1.22 M) + interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
+MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
 LIIF [11] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
@@
 Thera-L 1.41 M 34.88 31.22 29.24 26.96 23.89 22.33 21.32 20.61
-RDN [48] (22.0 M) MetaSR [17] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
+RDN [52] (22.0 M) + interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
+MetaSR [17] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
 LIIF [11] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
@@
 Thera-L 1.41 M 34.98 31.40 29.45 27.16 24.06 22.49 21.45 20.72
-SwinIR [23] (12.0 M) MetaSR [17] 0.45 M 35.15 31.40 29.33 26.94 23.80 22.26 21.26 20.54
+SwinIR [23] (12.0 M) + interpol. 35.09 31.42 29.48 27.17 24.08 22.53 21.52 20.82
+MetaSR [17] 0.45 M 35.15 31.40 29.33 26.94 23.80 22.26 21.26 20.54
 LIIF [11] 0.35 M 35.17 31.46 29.46 27.15 24.02 22.43 21.40 20.67
@@
 Thera-L 1.41 M 35.29 31.59 29.61 27.30 24.17 22.59 21.54 20.79
-Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the DIV2K validation set.
-The highest PSNR value per backbone and scaling factor is in bold and the second highest is underlined.
-We evaluate the performance of Thera on a variety of datasets and with three popular feature extractors as backbones.
-We also conduct an ablation study that supports the design choices described in Sec.
-3.
-Throughout this section, we will evaluate three variants of our method which differ solely in size of the hypernetwork and field:
-  • Thera-S: We set the number of shared components to 32, and omit all but one layer of the hypernetwork, the layer that maps features to field parameters.
-This version adds only 8,256 parameters on top of the backbone.
-  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork (5 ConvNeXt layers), approximately matching the parameter count of medium-sized competing methods (MetaSR [17], LIIF [11], LTE [22] and CUF [41]).
-This amounts to 0.48 M additional parameters.
-  • Thera-L: We use 512 components and a larger hypernetwork (16 ConvNeXt layers), amounting to 1.41 M additional parameters. This version is comparable to CiaoSR [8] in terms of parameter count.
-Datasets and metrics.
-Like previous methods, our models are trained with the DIV2K [1] training set, which is composed of 800 high-resolution RGB images of diverse scenes.
-We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [47], BSDS100 [27] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [28] (high-resolution Japanese manga illustrations).
-Following prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets.
-Not all numbers could be computed for competing methods due to the lack of availability of code or checkpoints, see Appendix E.
-Additional quantitative metrics are also provided in the supplementary material.
-Backbones.
-Following related work, we train our method jointly with three standard backbone feature extractors for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [48] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
 4.1 Super-Resolution Performance
-We begin with evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above.
+We start by evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above.
 Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors.
 Thera achieves state-of-the-art performance.
-Thera-L outperforms all previous methods on 22 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline).
-Thera-M, which requires only a third as many additional parameters, still outperforms or – less frequently – is on par with competitors that have comparable parameter count in most cases.
-Notably, except for ×$\times 2$, Thera-M also matches the performance of more expensive, attention-based methods CiaoSR and CLIT, despite a more than 30×30$\times$30 × fewer parameters than CLIT.
-Notably, our minimal version Thera-S – enabling arbitrary-scale SR with theoretical guarantees using only 8,192 additional parameters – regularly outperforms all comparison methods except for the most parameter-intensive ones when using the EDSR backbone.
-Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods (including attention-based ones) on 6 of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (and 170×170$\times$170 × fewer than CiaoSR) on top of the feature extractor.
+Thera-L outperforms all previous methods on 21 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline).
+Thera-M, which adds 3$\times$ fewer parameters to the backbone, still performs on par with or better than competitors that have comparable parameter counts.
+For scaling factors higher than ×$\times 2$, Thera-M also matches the performance of the more expensive, attention-based methods CiaoSR and CLIT, despite requiring 30×30$\times$30 × fewer parameters than CLIT.
+Notably, even our minimal version Thera-S, with the same anti-aliasing guarantees but with only 8,256 parameters on top of the backbone, reaches very competitive performance.
+It mostly outperforms all but the most parameter-intensive methods with the EDSR backbone.
+Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods, including attention-based ones, on 6 out of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (respectively 170×170$\times$170 × fewer than CiaoSR).
 Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors.
-This confirms that having formal modeling of multi-scale representation baked into our network is highly beneficial, especially for higher scaling factors.
+We see this as rather strong evidence for the power of the principled multi-scale observation model that is baked into our network; especially for higher scaling factors.
+For a comparison with conventional interpolation, we also report numbers using Lanczos (sinc) resampling with the respective ×$\times 4$ backbone.
+This baseline is almost always outperformed by dedicated ASSR methods, indicating that the latter do learn scale-specific priors.
+Table 2: Evaluation on common benchmark datasets for in-distribution scale factors, with a SwinIR backbone.
+The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
 Method Set5 Set14 B100 Urban100 Manga109
@@
 Thera-L 38.41 34.93 32.82 34.33 30.87 29.11 32.49 29.42 27.91 33.67 29.62 27.33 39.71 35.01 32.07
-Table 2: Evaluation on common benchmark datasets for in-distribution scale factors, with a SwinIR backbone.
-The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
-We further report the performance of our method on the five considered benchmark datasets in Tab.
+We further report the performance of Thera on five popular benchmark datasets in Tab.
 2.
 For these experiments, we use the SwinIR backbone.
-Overall, due to the out-of-distribution nature of the data and the small size of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent.
-This suggests that future work on ASSR would benefit from using larger-scale datasets in the future.
+Overall, due to the unreasonably small data volume and out-of-distribution nature of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent.
+We posit that future work on ASSR would benefit from switching to a larger, statistically more meaningful test bed.
 Nonetheless, our methods perform best in 12 out of 15 cases.
-We once again observe that the performance of Thera-S is often comparable to that of methods that require orders of magnitude more parameters.
-In our qualitative analysis – an excerpt of which is given in Fig.
-5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, in particular with repeating structure.
-This once again highlights the advantages of using an accurate observation model in our field formulation.
+We once again observe that the performance of Thera-S is often comparable to that of methods with orders of magnitude higher parameter counts.
+When visually inspecting the super-resolution results – for examples see Fig.
+5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, particularly in the presence of repeating structures.
+This again highlights the advantage of hard-wiring a theoretically principled sampling model that rules out signal aliasing.
+For further qualitative results see Appendix 0.B.
 DIV2K
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
-Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution. We use the SwinIR [23] backbone for all methods. More examples are given in Appendix B.
+LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
+Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution.
+We use the SwinIR [23] backbone for all methods.
+Differences between methods are highlighted with bounding boxes.
+More examples are given in Appendix 0.B.
 4.2 Multi-Scale Consistency
 We are interested in learning a super-resolution model that is consistent across scales.
-For instance, if we super-resolve an image with a scaling factor of $\times$4 then downsample it by a factor of $\times$2, we expect the result to be very similar to what we get by simply super-resolving the original image by a factor of $\times$2.
+For instance, if we super-resolve an image with a scaling factor $\times$4 then downsample it by a factor $\times$2, we expect the result to match what we get by directly super-resolving the original image by a factor $\times$2.
 We report in Tab.
-3 the quality of this match for several scaling factor pairs for Thera – which uses a formal observation model – and for three comparison methods that learn such model from data.
-First, we notice that better super-resolution quality does not correlate with scale consistency – e.g., LTE achieves better consistency than CiaoSR for most scaling factor pairs.
-Second, while other methods generally exhibit good consistency, even for out of distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
+3 the quality of this match for several pairs of scaling factors, both for Thera with its in-built observation model and for three comparison methods where the observation model must be discovered from data.
+First, we notice that higher super-resolution quality does not correlate with better scale consistency, e.g., LTE in most cases achieves better consistency than CiaoSR.
+Second, while other methods generally exhibit good consistency, even for out-of-distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
+Table 3: Multi-scale consistency (in PSNR) of super-resolution methods across various pairs of scaling factors.
 $(2,4)$ $(2,6)$ $(2,12)$ $(2,24)$ $(3,12)$ $(3,24)$ $(4,24)$
@@
 Thera-L 45.42 46.24 46.43 46.47 51.56 51.77 54.33
-Table 3: Comparison of multi-scale consistency (in PSNR) for various combinations of scaling factors.
 4.3 Ablation Studies
 In Tab.
-4, we ablate individual components and design choices of our method to understand their contribution to the overall performance.
-These comparisons are done using Thera-M using the EDSR-baseline backbone.
+4, we ablate individual components and design choices of our method to understand their contributions to overall performance.
+These comparisons use Thera-M with EDSR-baseline backbone, but are also representative of other settings.
 Single scale training.
-We ran three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization.
-$\kappa$ was fixed for these experiments since it requires multiscale training to be optimized.
-We observe equal or even superior performance compared to our base case for the scale used for training (marked in yellow in Tab.
-4), but we see a significant drop in generalization.
-Trainable $\kappa$.
-We observe a small drop in performance when using the fixed, theoretically derived value for $\kappa$ (Eq.
-5).
-This suggests that there are still effects that have not been accounted for in our model, but the effects are very minor.
+We ran three experiments using a single scale ($\times$2, $\times$3, or $\times$4) to test how this affects scale generalization.
+$\kappa$ was fixed for these experiments as multi-scale training is required to optimize it.
+As expected, we observe equal or even superior performance compared of single-scale training when tested at the training scale (marked in yellow in Tab.
+4), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
+Trainable $\bm{\kappa}$.
+Fixing $\kappa$ at the theoretically derived value (Eq.
+5) leads to a small drop in performance.
+This suggests that there remain effects that are not accounted for by our proposed observation model, but that these are very minor.
 Geometric self-ensemble.
-As observed by previous work [40, 24, 41], we see a notable boost in performance when using geometric self-ensemble (GSE).
-Nevertheless, if a user prioritizes speed over quality, this addition can be disabled without retraining the network at test-time.
-Total variation. We observe negligible effects for in-domain scaling factors when no TV is used, but performance degrades significantly for out-of-distribution scales.
+In line with previous work [42, 24, 43] we see a notable performance boost with geometric self-ensembling (GSE).
+Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
+Total variation. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
 No thermal activations.
 We replace thermal activations (Eq.
-4) for standard ReLU activations.
+4) with standard ReLU activations.
 What remains is only the hypernetwork controlling the parameters of the local fields.
-A loss in performance shows the value of our proposed thermal activation for multiscale representation.
+A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
 Shared components.
-We observe a noticeable performance drop in all scales if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$.
-Despite being even more flexible, it seems challenging to learn per-field components and other correlated parameters simultaneously.
+We observe a noticeable performance drop, at all scales, if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$.
+Despite the greater flexibility afforded by individual per-field components, it seems challenging to learn them jointly with other, correlated parameters.
+Table 4: Ablation experiments with Thera-M using the EDSR-baseline backbone.
 Method In-distribution Out-of-distribution
@@
 No TV prior 34.87 31.21 29.21 26.86 20.82 19.12
-No $\xi$ 34.80 31.12 29.13 26.87 22.28 20.58
+ReLU instead of $\xi$ 34.80 31.12 29.13 26.87 22.28 20.58
 Non-shared comps. 34.80 31.13 29.14 26.87 22.28 20.57
-Table 4: Ablations have been run using Thera-M using the EDSR-baseline backbone, a representative case for all other settings.
 4.4 Limitations and Future Work
@@
 While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder.
-While we have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work.
-This could result in higher parameter efficiency, and potentially better generalization.
-There may also be other formulations with similar characteristics which do not require the usage of a hypernetwork, which may allow for more streamlined training.
-We noticed a relative drop in performance w.r.t.
-other methods for $\times$2 SR.
-We speculate that our weakness relative to higher scaling factors is due to this scale being close to the interface between continuous (field) and grid (pixel) representations, which Thera’s formalism does not model.
-Further exploration on this transition regime for such hybrid structures would be useful to better understand and use these hybrid representations.
-This could also benefit Instant-NGP [31] and related methods.
-Finally, we have only explored TV as a method for regularizing $\Phi$ at $t=0$, but we expect other signal priors to be even more effective.
-Priors at $t=0$, possible only by using Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.
+We have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, and we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work.
+This could result in even higher parameter efficiency, and potentially better generalization.
+It may also be possible to construct similar formulations that do not require the usage of a hypernetwork, which could enable more streamlined training.
+Relative to other methods, Thera’s performance slightly drops for $\times$2 super-resolution when compared to higher scaling factors.
+We speculate that at small scaling factors Thera’s isotropic, continuous field representation is increasingly challenged by anisotropic effects from the oriented, discrete pixel grid.
+Further explorations are necessary to better understand this transition regime and perhaps come up with hybrid representations (which could potentially benefit also other methods at the interface of gridded and implicit representations like Instant-NGP [33]).
+Finally, we expect that more advanced priors than TV could be even more effective at regularizing $\Phi$.
+Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.
 5 Conclusion
 We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations.
-Our method, Thera, implicitly describes an image as a combination of sinusoidal components which can be modulated based on their frequency to perform Gaussian filtering (anti-aliasing) analytically with little overhead.
-Our experimental evaluation shows that Thera is superior to competing methods in most cases, while also being more parameter efficient and providing theoretical guarantees.
-We also hope to inspire others to apply Thera to other problems where it is likely to be beneficial due to its Gaussian observation model, such as demosaicing, burst super-resolution, and neural radiance fields.
-Supplementary Material
-Appendix A Theory
-A.1 Preliminaries
+Our method, Thera, implicitly describes an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) analytically, with negligible overhead.
+Our experimental evaluation shows that Thera outperforms competing methods in most cases, while at the same time being more parameter efficient, and offering theoretical guarantees.
+We are convinced that Thera-style representations could also benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful observation models like our Gaussian PSFs.
+Appendix 0.A Theory
+0.A.1 Preliminaries
 As was described in Sec.
-3.1, the idea is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
+3.1, the idea underlying our neural field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
 $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;.$ (8)
-The reason for this is that the analytical solution to the (isotropic) heat equation can be modelled as a convolution of the initial state $\Phi(\mathbf{x},0)$ with a Gaussian kernel
+The reason for this is that the analytical solution to the (isotropic) heat equation can be modeled as a convolution of the initial state $\Phi(\mathbf{x},0)$ with a Gaussian kernel
 $g(\mathbf{x},t)=\frac{1}{4\pi\kappa t}\cdot\exp\left(-\frac{x_{1}^{2}+x_{2}^{2}}{4\kappa t}\right).$ (9)
-By fitting the data at $\Phi(\mathbf{x},1)$, we are assuming a Gaussian point spread function (PSF) with the shape
+By fitting the data (image $I$) at $\Phi(\mathbf{x},1)$, we are assuming a Gaussian point spread function (PSF) with the shape
 $\text{PSF}(\mathbf{x})=\frac{1}{4\pi\kappa}\cdot\exp\left(-\frac{x_{1}^{2}+x_{2}^{2}}{4\kappa}\right).$ (10)
@@
 $\Phi(\mathbf{x},t)\text{ is}\begin{cases}\text{meaningless},&\text{if}\ t<0\\ \text{pure signal},&\text{if}\ t=0\\ \text{ill-posed},&\text{if}\ 0<t<1\\ I,&\text{if}\ t=1\\ \text{well-posed},&\text{if}\ t>1\\ \end{cases}$ (11)
-The ill-posed problem for $10<t<$ 1 is the interesting case. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
+The ill-posed problem for $10<t<$ 1 is the interesting case, where this formulation relates to super-resolution.
+The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
 For all the formulations here, we define the image $I$ to correspond to the coordinates $x_{1},x_{2}\in[-0.5,0.5]$.
-A.2 Thermal diffusivity coefficient
+0.A.2 Thermal Diffusivity Coefficient
 To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$.
 One way to do so is to match the cut-off frequency of the filter in Eq.
-9 at $t=1$ to the well-known Nyquist frequency given by the sampling rate.
+9 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate.
 We take the cut-off frequency of the Gaussian filter defined in Eq.
@@
 $f_{\text{Nyquist}}=\frac{N}{2},$ (13)
-where $N$ is the number of samples along a given dimension. This formulation assumes even sampling over $x_{1}$ and $x_{2}$.
+where $N$ is the number of samples along a given dimension.
+This formulation assumes even sampling over $x_{1}$ and $x_{2}$.
+To extend this formulation to non-square images, it would be necessary to change the shape of the signal’s domain in order to maintain even sampling in all spatial dimensions.
 If we solve for $f_{c}=f_{\text{Nyquist}}$ at $t=1$, we get
 $\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}N^{2}}.$ (14)
-For the Thera formulation presented in Sec. 3.2, we want $\Phi$ to contain a single pixel at $t=1$, therefore we initialize $\kappa$ with
+For the Thera formulation presented in Sec.
+3.2, we want $\Phi$ to contain a single pixel at $t=1$.
+This is the pixel from the low-resolution input which will become $SR^{2}$ pixels for super-resolution with a scaling factor of $SR$.
+Therefore we initialize $\kappa$ with
 $\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}}.$ (15)
-Note that the exact value of $\kappa$ will depend on the exact characteristics of the system that is being modelled.
-Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but is also more prone to aliasing.
+Note that the exact value of $\kappa$ will depend on the characteristics of the system that is being modeled and the anti-aliasing filter that was used (or is assumed).
+Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but are also more prone to aliasing.
 Finally, we would like to highlight that Eq.
 12 is specific to the case where $\mathbf{x}$ is 2-dimensional.
-The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using fields in spaces with numbers of spatial dimensions other than 2.
-Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the desired formulas for a Gaussian filter with the appropriate number of dimensions.
-A.3 Relationship between $t$ and $s$
+The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using our field’s formulation in spaces with numbers of spatial dimensions other than 2.
+Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the formulas for a Gaussian filter with the appropriate number of dimensions.
+0.A.3 Relationship Between $t$ and $s$
 Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate.
-If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it with $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$.
+If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it down to $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$.
 Using Eq.
@@
 $t=0.5^{2}=0.25.$ (17)
-Thus, $10<t<$ 1 refer to super-resolution, while $t>1$ refers to downsampling.
+Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling.
+This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
+0.A.4 Other Filters
+In theory, the formulation presented in Sec.
+3.1 allows us to use any low-pass filter we want, since we can modulate different components freely.
+Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation.
+Initial explorations of a sharp low-pass filter that completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely due to the associated effect on gradients during training.
+It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way.
+For quantitative evaluations, this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios, this may be desirable.
+0.A.5 Initialization of Components
+We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Eq.
+2, is important.
+Sitzmann et al.
+[40] made similar observations when periodic activation functions were first used for neural fields.
+The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training.
+Thus, we choose to initialize $\mathbf{W}_{1}$ such that
+$p(\left|\nu(\mathbf{w}_{1})\right|)\propto\left|\nu(\mathbf{w}_{1})\right|$ (18)
+up to a given maximum frequency, allotting more components to higher frequencies. See code for more details.
+Appendix 0.B Additional Qualitative Results
+Figures 9 and 10 contain additional qualitative results for lower and higher scaling factors, respectively.
+Again, we see that Thera can reconstruct the targets more faithfully.
+This is particularly (but not exclusively) visible in repetitive structural elements.
+It seems evident that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this.
+Furthermore, we provide a showcase of the continuous upsampling capabilities of our method in Fig.
+6.
+Figure 6: Showcase of multiscale upsampling using Thera-L with a SwinIR [23] backbone, shown with non-integer scaling factors.
+Appendix 0.C Additional Quantitative Results
+Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab.
+1.
+We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric.
+Furthermore, Tab.
+6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
+Finally, in Fig.
+7 we compare the number of additional parameters and PSNR values for various methods and upsampling factors on DIV2K.
+Figure 7: Plots depicting the number of additional parameters vs.
+PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 2$ (left), ×$\times 3$ (center) and ×$\times 4$ (right) upsampling tasks.
+Table 5: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the SwinIR [23] backbone for all models.
 In-distribution Out-of-distribution
@@
 Thera-L .942 .885 .834 .759 .708 .646 .597 .570 .552
-Table 5: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the SwinIR [23] backbone for all models.
+Table 6: Evaluation on common benchmark datasets for out-of-distribution scale factors, using the SwinIR backbone.
+The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
 Method Set5 Set14 B100 Urban100 Manga109
@@
 Thera-L (ours) 29.48 27.46 25.09 26.94 25.45 23.42 26.14 25.08 23.71 24.79 23.32 21.62 28.13 25.69 22.85
-Table 6: Evaluation on common benchmark datasets for out-of-distribution scale factors, using the SwinIR backbone.
-The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
-A.4 Other filters
-In theory, the formulation presented in Sec.
-3.1 allows us to use any low-pass filter we want, since we can modulate different components freely.
-Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation.
-Initial explorations of a sharp low-pass filter which completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely to how this also affected gradients during training.
-It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way.
-For quantitative evaluations this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios this may be desirable.
-A.5 Initialization of components
-We have noticed during our experiments that the initialization of the components ($\mathbf{W}_{1}$ in Eq.
-2) is very important.
-Sitzmann et al.
-[38] made similar observations when periodic activation functions were first used for neural fields.
-The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training.
-Thus, we chose to initialize $\mathbf{W}_{1}$ such that
-$p(\left|\nu(\mathbf{w}_{1})\right|)\propto\left|\nu(\mathbf{w}_{1})\right|$ (18)
-up to a given maximum frequency, allotting more components to higher frequencies. See code for more details.
-Appendix B Additional qualitative results
-Figure 8 and Fig.
-9 contain additional qualitative results for lower and higher scaling factors, respectively.
-Again, we see that Thera can reconstruct the targets more faithfully.
-This is particularly visible in repetitive structural elements.
-We suspect that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this.
-Appendix C Additional quantitative results
-Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab.
-1.
-We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric.
-Furthermore, Tab.
-6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
-Finally, in Fig.
-7 we show the number of additional parameters for various methods on the ×$\times 3$ upsampling task.
-Appendix D Hypernetwork architecture
-Figure 6: Schematics of the hypernetworks for a) Thera-S, b) Thera-M and c) Thera-L.
+Appendix 0.D Hypernetwork Architecture
 In Fig.
-6, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L.
+8, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L.
 The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image.
 It is then followed by a number of (residual) ConvNeXt [26] blocks.
-Each block consists of i) a $3\times 3$ convolution that maps to $d$-dimensional features, ii) a layer normalization [2] operation, iii) a $1\times 1$ convolution with $4d$ filters, iv) a GELU [15] non-linearity, and v) another $1\times 1$ convolution with $d$ filters.
+Each block consists of (i) a $3\times 3$ convolution that maps to $d$-dimensional features, (ii) a layer normalization [2] operation, (iii) a $1\times 1$ convolution with $4d$ filters, (iv) a GELU [15] non-linearity, and (v) another $1\times 1$ convolution with $d$ filters.
 Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al.
@@
 For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
+Figure 8: Schematics of the hypernetworks for a) Thera-S, b) Thera-M and c) Thera-L.
 Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [40].
 There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L.
-Let $c$ be the number of components used (32, 512 and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork.
-The total parameter count of the hypernetwork is 8,192, 0.48 M and 1.41 M for the three variants, respectively.
+Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork.
+The total parameter count of the hypernetwork is 8,192, 0.48 M, and 1.41 M for the three variants, respectively.
 To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters.
 We highlight that these do not come from the hypernetwork.
-Appendix E Replication of other methods
-We encountered challenges attempting to recreate results of some of the competitor methods, which explains why some numbers are missing in various tables or differ from the originally reported ones.
+Appendix 0.E Replication of Other Methods
+We encountered challenges attempting to recreate results of some of the competing methods, which explains why some numbers are missing in various tables or differ from the originally reported ones.
 Details are provided below.
 CUF [43].
-At the time of writing, there was no public code or checkpoints available for CUF.
+At the time of writing, there were no public code or checkpoints available for CUF.
 Therefore, we could not generate numbers for datasets and scaling factors not reported in the original paper.
@@
 For CLIT, at the time of writing, there is public code available, however, checkpoints are not.
-We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU hours: over 1 month using 8×8$\times$8 × Nvidia GeForce RTX 3090s.
+We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU time: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090s.
 The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
 CiaoSR [8].
-We noticed that CiaoSR’s border cropping prior to evaluation on the DIV2K dataset in the official implementation deviated slightly from the rest of the comparison methods.
-We aligned this to align the evaluation of all methods and re-computed all DIV2K numbers.
+We noticed that in the official CiaoSR implementation, border cropping prior to evaluation on the DIV2K dataset deviated slightly from all other methods.
+We aligned this to enable a meaningful comparison and re-computed all DIV2K numbers.
 This explains the slight deviation from the original paper.
-Figure 7: Plots depicting the number of additional parameters vs. PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 3$ upsampling task.
 DIV2K ×$\times 2$
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Urban100 ×$\times 3$
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 Manga109 ×$\times 3$
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
+LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
 Figure 9: Further qualitative examples for ×$\times 2$ and ×$\times 3$ upsampling factors. Best viewed zoomed in.
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
+LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
 Figure 10: Further qualitative examples for ×$\times 4$ and ×$\times 6$ upsampling factors. Best viewed zoomed in.
+Appendix 0.F Computational Requirements
+For comparison of computational requirements between methods, we provide training time, inference time, and inference VRAM requirements in Tab.
+7, for the EDSR-baseline backbone.
+These values assume NVIDIA GeForce RTX 3090 Ti GPUs, and they show that we are more efficient than recent transformer-based methods, but still reach superior performance, as shown in the main paper.
+Table 7: Computational requirements of various methods.
+w/ EDSR-b Training time Inference time VRAM
+(GPU days)
+MetaSR 0.42 3.68 ms 588 MiB
+LIIF 0.71 13.7 ms 1174 MiB
+LTE 0.79 17.1 ms 1250 MiB
+CiaoSR $\sim 2.6$ 37.7 ms 6474 MiB
+CLIT $\sim 4.6$ 16.0 ms 2984 MiB
+Thera-S 1.0 3.52 ms 754 MiB
+Thera-M 2.2 10.0 ms 2290 MiB
+Thera-L 3.3 14.3 ms 2292 MiB

### v3

@@
-Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution
+Thera: Aliasing-Free Arbitrary-Scale Super-Resolution with Neural Heat Fields
 Abstract
-Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at arbitrary rates.
-However, the point-wise query of the neural field does not naturally match the point spread function (PSF) of a given pixel, which may cause aliasing in the super-resolved image.
-We present a novel way to design neural fields such that points can be queried with an adaptive Gaussian PSF, so as to guarantee correct anti-aliasing at any desired output resolution.
-We achieve this with a novel activation function derived from Fourier theory.
-Querying points with a Gaussian PSF, compliant with sampling theory, does not incur any additional computational cost in our framework, unlike filtering in the image domain.
-With its theoretically guaranteed anti-aliasing, our method sets a new state of the art for ASSR, while being more parameter-efficient than previous methods.
-Notably, even a minimal version of our model still outperforms previous methods in most cases, while adding 2-4 orders of magnitude fewer parameters.
-Code and pretrained models are available at https://github.com/prs-eth/thera.
+Recent approaches to arbitrary-scale single image super-resolution (ASR) use neural fields to represent continuous signals that can be sampled at arbitrary resolutions.
+However, point-wise queries of neural fields do not naturally match the point spread function (PSF) of pixels, which may cause aliasing in the super-resolved image.
+Existing methods attempt to mitigate this by approximating an integral version of the field at each scaling factor, compromising both fidelity and generalization.
+In this work, we introduce neural heat fields, a novel neural field formulation that inherently models a physically exact PSF.
+Our formulation enables analytically correct anti-aliasing at any desired output resolution, and – unlike supersampling – at no additional cost.
+Building on this foundation, we propose Thera, an end-to-end ASR method that substantially outperforms existing approaches, while being more parameter-efficient and offering strong theoretical guarantees.
+The project page is at https://therasr.github.io.
+Figure 1: We present Thera, the first method for arbitrary-scale super-resolution with a built-in physical observation model.
+Given an input image, a hypernetwork predicts the parameters of a specially designed neural heat field, inherently decomposing the image into sinusoidal components.
+The field’s architecture automatically attenuates frequencies as a function of the scaling factor so as to match the output resolution at which the signal is re-sampled.
 1 Introduction
-For years, learning-based image super-resolution (SR) methods have achieved ever better results.
-But unlike interpolation methods, which allow us to resample a given image at any resolution, those methods require retraining for each desired scaling factor.
-Recently, several schemes for arbitrary-scale SR (ASSR) have been proposed that allow end-users to choose any scaling factor, rendering the method much more flexible [17].
+Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results.
+However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor.
+Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility [18].
 Notably, with LIIF, Chen et al.
 [10] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors.
-Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10].
+LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields [24, 7, 9, 56].
 This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [48].
-By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
-Among others, they have been widely adopted for 3D reconstruction with radiance [32], occupancy [31, 37], and distance fields [35, 39].
-Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components.
-We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies (insets illustrate components decaying at different rates).
-This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
-Neural fields offer a compelling representation of continuous signals.
-However, for many tasks, it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3].
-ASSR calls for a neighborhood-dependent observation model, as the signal should be observed at various sampling frequencies and is prone to aliasing.
+By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
+While neural fields naturally model continuous functions, they do not easily allow for observations of such functions other than point-wise evaluations.
+For many tasks, however, integral observation models such as point spread functions (PSFs) are desirable.
+This is particularly true for neural fields-based ASR methods, which by nature do not commit to a fixed upscaling factor a priori but regress continuous representations with unbounded spectra that can be observed at various sampling rates.
+If the Nyquist frequency corresponding to the desired sampling rate is lower than the highest frequency represented by the field, the sampling operation is prone to aliasing.
+This explains the initially counterintuitive relevance of anti-aliasing for super-resolution: When using neural fields, signals are first upsampled to infinite (continuous) resolution and then resampled at the desired resolution, and this latter operation must be done carefully.
 Incorporating a physically plausible observation model is not trivial [2, 3, 27, 51, 17, 4], but has the potential to avoid aliasing.
 For this reason, Chen et al.
-[11] already took a first step towards learning multiscale representations, via cell encoding.
-For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the input coordinates, as a means to control different frequencies in the field.
-Lindell et al.
-[25] and Yang et al.
-[49] have succeeded in performing subband decomposition of neural fields, though bands are architecturally fixed to a small number of discrete intervals and cannot be modified after training.
-In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction.
-Interestingly, this filtering incurs no additional overhead: The cost of querying field values is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
-The key insight is that sinusoidal activation functions [40] offer the possibility to selectively decay each component according to its spatial frequency, following Fourier theory.
-Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, a neural field method for ASSR underpinned by a novel thermal activation function, see Fig.
-1.
-That formulation makes it possible to analytically compute Gaussian-blurred versions of the neural field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius.
-I.e., one can query $\Phi$ with a Gaussian PSF that is matched to the output resolution to prevent aliasing.
-[^1] While our framework is in principle not bound to the heat equation or Gaussian PSFs, it is an intuitive interpretation for readers familiar with signal theory.
-Thera provides theoretical guarantees w.r.t.
-anti-aliasing and multi-scale representation, in exchange for certain constraints on the neural architecture that can easily be overcome with a hypernetwork [14].
-Empirically, it outperforms competing methods in most of our experiments and is far more parameter-efficient.
-We demonstrate that a lightweight variant of our model still outperforms most previous methods while adding 2-4 orders of magnitude fewer parameters.
-To the best of our knowledge, Thera is the first neural field method to allow bandwidth control at test time.
-Our main contributions are:
+[10] and successor works [24, 7, 9, 56] have already taken a first step towards learning multi-scale representations, via cell encoding.
+Fundamentally, these “learning-based anti-aliasing” approaches try to learn an integrated version of the field for any scaling factor, wasting field capacity for approximating a relation that can be described exactly through Fourier theory.
+$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$31.531.531.531.531.631.631.631.631.731.731.731.731.831.831.831.831.931.931.931.93232323232.132.132.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
+Figure 2: Comparison of recent ASR methods, averaged over ×$\times\{2,3,4\}$ scales.
+We generally achieve higher performance at lower parameter counts.
+Our best model, Thera Pro, achieves highest overall performance by a large margin.
+In this work, we combine recent advances in implicit neural representations with ideas from classical signal theory to introduce neural heat fields, a novel type of neural field that guarantees anti-aliasing by construction.
+The key insight is that sinusoidal activation functions [41] enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory.
+This allows for the exact computation of Gaussian-blurred versions of the field for any desired (isotropic) blur radius.
+When rasterizing an image, the field can therefore be queried with a Gaussian PSF that matches the target resolution, effectively preventing aliasing.
+Notably, filtering with neural heat fields incurs no computational overhead: The querying cost is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
+Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork [16] with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Fig.
+1).
+Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Fig.
+2).
+To the best of our knowledge, Thera is also the first neural field method to allow bandwidth control at test time.
+In summary, our main contributions are:
   1. 1.
-    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and therefore anti-aliasing with minimal overhead.
+    We introduce neural heat fields, which represent a signal with a built-in, principled Gaussian observation model, and therefore allow anti-aliasing with minimal overhead.
   2. 2.
-    We use thermal activations to build Thera, a novel method for ASSR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
+    We use neural heat fields to build Thera, a novel method for ASR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
 2 Related Work
-2.0.1 Arbitrary-scale SR.
-In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods.
+2.1 Neural Fields
+A neural field, also called an implicit neural representation, is a neural network trained to map coordinates onto values of some physical quantity.
+Recently, neural fields have been used for parameterizing various types of visual data, including images [22, 41, 42, 10, 24, 12, 47], 3D scenes (e.g., represented as signed distance fields [36, 41, 40, 46, 47], occupancy fields [33, 37], LiDAR fields [20], view-dependent radiance fields [34, 2, 3, 4, 47]), or digital humans [52, 55, 8, 49, 15].
+Frequently, it is desirable to impose some prior over the space of learnable implicit representations.
+A common approach for such conditioning is encoder-based inference [48], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [10, 24, 44, 7, 9].
+The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [16], mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$.
+An early example of this approach, which is gaining popularity [48], was proposed in [41].
+2.2 Arbitrary-Scale Super-Resolution
+ASR is the sub-field of single-image SR in which the desired SR scaling factor can be chosen at inference time to be (theoretically) any positive number, allowing maximum flexibility, such as that of interpolation methods.
 The first work along this line is MetaSR [18], which infers the parameters of a convolutional upsampling layer using a hypernetwork [16] conditioned on the desired scaling factor.
-An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by local neural fields that are conditioned with features extracted from the low-resolution input image.
-The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution.
-Following work has since built upon the LIIF framework.
-For instance, UltraSR [48] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields [32, 3, 4, 5].
+An influential successor work is LIIF [10], in which the high-resolution image is implicitly described by local neural fields.
+These fields are conditioned via concatenation, with features extracted from the low-resolution input image.
+The continuous nature of the neural fields allows for sampling target pixels at arbitrary locations and thus also arbitrary resolution.
+Most subsequent work has since been built upon the LIIF framework.
+For example, UltraSR [50] improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields [34, 2, 3, 4].
 LTE [24] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP.
@@
 [44] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling.
-Most recently, methods have integrated attention mechanisms into their implicit representation, enabling the neural field to adaptively attend to image features depending on the query coordinates and scale factor, improving reconstruction quality [8, 10].
-These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data.
-In this paper, we explore a way to directly integrate a (Gaussian) observation model into the neural field representation.
-Implicit neural representations.
-A neural field is a neural network trained to map coordinates onto values of some physical quantity.
-Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 40, 41, 11, 22, 12, 45], 3D scenes (e.g., represented as signed distance fields [35, 40, 39, 44, 45], occupancy fields [31, 37], LiDAR fields [19], view-dependent radiance fields [32, 3, 4, 5, 45]), or digital humans [50, 53, 9, 47, 13].
-We refer the interested reader to Xie et al.
-[46] for a comprehensive survey of neural fields in vision and beyond.
-Frequently, it is desirable to impose some prior over the space of learnable representations.
-A common approach for such conditioning is encoder-based inference [46], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 43, 8, 10].
-The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [14].
-The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, directly mapping latent codes to parameters of the field $\bm{\theta}$.
-An early example of this approach, which is gaining popularity [46], was proposed by Sitzmann et al.
-[40].
-Anti-aliasing in neural fields.
+More recently, methods like CiaoSR [7], CLIT [9], and most recently MSIT [56] have integrated (multi-scale) attention mechanisms, improving reconstruction quality.
+In a parallel line of research, Wei et al.
+[45] propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
+2.3 Anti-Aliasing in Neural Fields
 Early in the recent development of implicit neural representations, concerns regarding aliasing were raised.
 Barron et al.
-[3] proposed integrating the positional encoding with Gaussian weights, which reduced aliasing in NeRF.
+[2] proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF [34].
 Improvements were later proposed for unbounded scenes [3] and to improve efficiency [17].
-Barron et al.
-[5] tackle anti-aliasing for the increasingly popular Instant-NGP [33] approach.
-Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [25] or polynomial neural fields [49], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks.
-Such representations would not be a good fit for ASSR since they do not allow for continuous anti-aliasing, nor bandwidth control at test time.
-In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
+In [4], Barron et al.
+tackle anti-aliasing within the Instant-NGP [35] approach.
+Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [27], polynomial neural fields [51] or cascaded training [39], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks.
+These methods are not a good fit for ASR because they do not allow for continuous anti-aliasing, nor bandwidth control at test time.
+To perform scale-dependent filtering, most fields-based ASR methods instead explicitly provide the scale as input to the field, attempting to learn an appropriate observation model from data.
+While this approach may work reasonably well in in-distribution settings, it seeks to learn a model from data that can be described exactly with a differential equation, ultimately sacrificing fidelity and generalization.
+In contrast, in this paper we explore a way to directly integrate a physics-informed observation model into the neural field representation.
 3 Method
-To expound how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis.
-For simplicity, we stick to 2-dimensional signals $\mathbf{x}\in\mathbb{R}^{2}$, but all theory is valid for any number of dimensions.
-We refer the reader to [34] for a comprehensive discussion of sampling and aliasing.
-To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate.
-One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$.
-We will show that one can naturally integrate this behavior into a neural field’s architecture.
-The key to this is that, if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor.
-This equivalence is depicted in Fig.
-2.
-Decomposing a given signal into sinusoidal components is normally a costly operation, but it turns out that we can avoid it by designing the neural field such that the decomposition itself is the native representation of the signal.
-3.1 Thermal Activation Functions
-The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation.
-Other implicit representation methods, such as SIREN [40], optimize the parameters $\bm{\theta}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$.
-Our key insight is that we can interpret this field as the temperature distribution on an (isotropic) surface at a given time $t_{0}$.
-It is well known how to solve the heat equation forwards in time – its solution is obtained by blurring the initial state at $t_{0}$ with a Gaussian kernel whose radius depends on the time difference $t-t_{0}$.
-Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
-$\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;,$ (1)
-with $\kappa>0$ the thermal diffusivity of the medium.
-In contrast, the solution for $0\leq t<t_{0}$ is ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
-Figure 2: If we have access to the sinusoidal decomposition of a signal, filtering can be done analytically by scaling each component appropriately.
-We design our neural field so that we have access to such a decomposition at no extra cost.
-It turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
+Figure 3: Overview of Thera.
+A hypernetwork estimates parameters $\{\bm{b}_{1},\bm{W}_{2}\}^{(i,j)}$ of pixel-wise, local neural heat fields.
+The phase shifts $\bm{b}_{1}$ operate on globally learned components, before thermal activations scale each component depending on their frequency and the desired scaling factor.
+The components are then linearly combined using coefficients $\bm{W}_{2}$, resulting in an appropriately-blurred, continuous local neural field.
+This field is then rasterized at the appropriate sampling rate (resolution) to yield a part of the final output image (red square).
+In this section we introduce Thera, a novel neural fields-based ASR method that guarantees analytical anti-aliasing at any desired output resolution at no additional cost.
+First, we present neural heat fields, a special type of neural field that inherently achieves anti-aliasing by implicitly attenuating high-frequency components as a function of a time coordinate.
+Next, we propose a mechanism for learning a prior over a grid of neural heat fields, enabling them to represent a multi-scale output image conditioned on a lower-resolution input image.
+Finally, we show that our formulation allows us to impose a regularizer on the underlying, continuous signal itself – something that, to the best of our knowledge is not possible in previous methods.
+3.1 Neural Heat Fields for Analytical Anti-Aliasing
+Let $\mathbf{x}\in\mathbb{R}^{2}$ denote the spatial coordinates of a continuous image function $f(\mathbf{x})$.
+Aliasing occurs when this continuous signal is sampled at a rate that does not adequately capture its highest frequency components, resulting in overlapping spectral replicas in the Fourier domain.
+One must therefore apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$.
+The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency and a time coordinate $t$.
+This behavior is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
 $\Phi(\mathbf{x},t)$ $=\mathbf{W}_{2}\cdot\xi\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t\right)+\textbf{b}_{2},$ (1)
-where $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$ are MLP parameters, $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$ (described in more detail later), and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time.
-We set $t=0$ to be the initial thermal state of a given system represented as a sum of sinusoidal components:
-$\xi(\mathbf{z},\nu,\kappa,0)=\sin(\mathbf{z}).$ (3)
-For $t>0$, each component is decayed according to its wave number, the thermal diffusivity constant $\kappa$, and $t$:
+with parameters $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$.
+Intuitively, $\mathbf{W}_{1}$ serves as a frequency bank, with its components acting as the basis functions that compose the signal $\Phi(\mathbf{x},0)$, and phase shifts encoded by $\mathbf{b}_{1}$.
+The matrix $\mathbf{W}_{2}$, with one row per output channel, contains initial magnitudes of these components, and $\mathbf{b}_{2}$ is the global bias of $\Phi$ per channel.
+Finally, we introduce the thermal activation function $\xi(\cdot)$, which models the aforementioned decay of sinusoidal components (implied by $\mathbf{W}_{1}$) over time:
 $\xi(\mathbf{z},\nu,\kappa,t)=\sin(\mathbf{z})\cdot\exp(-\left|\nu\right|^{2}\kappa t).$ (2)
-To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally, regardless of their orientation).
-Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
-There is an ideal bijection between the sampling rate $f_{s}$ and $t$.
+Here, $\left|\nu\right|=\left|\nu(\mathbf{W}_{1})\right|$ denotes the row-wise Euclidean norm of $\mathbf{W}_{1}$, representing the magnitudes of the implied wave numbers (frequencies).
+Interestingly, Eq.
+1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A.
+We therefore refer to this MLP as a neural heat field.
+There is an ideal bijection between the desired sampling rate $f_{s}$ and $t$.
 At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$).
@@
 To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$.
-The relationship between the cut-off frequency of the filter and $t$ is controlled by $\kappa$, which we can freely set to any positive number.
+The relationship between the cut-off frequency of the filter and $t$ is controlled by the diffusivity constant $\kappa$, which we can freely set to any positive number.
 For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$.
 Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
-$\kappa=\frac{\sqrt{\log(4)}}{2f_{s}\pi^{2}}.$ (5)
-To subsample the signal $\mathcal{D}$ by a factor $S$, $\Phi$ should be sampled at
+$\kappa=\frac{\sqrt{\ln(4)}}{2f_{s}\pi^{2}}.$ (3)
+To subsample the signal $\mathcal{D}$ by a factor $S$, the field $\Phi$ should be sampled at
 $t=S^{2}.$ (4)
-The derivations of these values can be found in Appendix 0.A.
-In our formulation, $\mathbf{W}_{1}$ can be seen as a frequency bank whose components (“basis functions”) make up the signal $\Phi(\mathbf{x},0)$, and the vector $\mathbf{b}_{1}$ represents the phase shifts of those components.
-The matrix $\mathbf{W}_{2}$, with one row per output channel, contains the initial ($t=0$) magnitudes of these components, and $\mathbf{b}_{2}$ is the global bias of $\Phi$ per channel.
-Image domain
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA $\times$1/32 AA
-Neural field
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$ $\Phi(\mathbf{x},32^{2})$
-Figure 3: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with thermal activations (bottom row), which is computed without convolutions or over-sampling.
-The neural field was supervised only at $\Phi(\mathbf{x},1)$.
-Figure 3 shows an example where we fit the field at $t=1$ to the image.
-After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq.
-6.
-We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
-By definition, our field $\Phi(\mathbf{x},t)$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies.
-Note that modulating deeper fields, as is done in [30], would no longer guarantee anti-aliasing.
-An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.
-3.2 Thermal Activations for Super-Resolution
-The multi-scale signal representation defined above is a natural match for super-resolution, and particularly well-suited if arbitrary scale factors shall be covered.
+For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Appendix A.
+3.2 Learning a Super-Resolution Prior
+The multi-scale signal representation inherent in neural heat fields is a natural match for ASR.
 Still, two challenges must be addressed.
 First, our formulation restricts the choice of architecture to MLPs with a single hidden layer.
-Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed.
+Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($10<t<$ 1) remains ill-posed.
 To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data.
-Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size.
-The hypernetwork maps the low-resolution input image to parameters of local fields, each spanning the area of one pixel of the low-resolution input as originally proposed by LIIF [11] and adopted by most recent ASSR methods [22, 8, 43, 10].
-Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field.
+Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, mapping the low-resolution input image to the $N$ parameters of each local neural heat field.
+As originally proposed by LIIF [10] and adopted by recent ASR methods [24, 7, 44, 9], each local field spans the area of one pixel of the low-resolution input.
+It is important that even though the fields themselves model only a local part of the image, the hypernetwork informs them with contextual features collected over a large receptive field.
 The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [10, 24, 44, 7, 9], followed by one or more layers that produce feature maps of depth $N$.
-In this work, we use a series of ConvNeXt [26] blocks.
-The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$.
-The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values.
-Similar to LIIF, the local coordinate system for every field is bounded by $[-0.5,0.5]\times[-0.5,0.5]$.
+During training, the local fields are supervised with values of high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$ (ensuring that the signal is correctly blurred for the target resolution), and the entire architecture is optimized end-to-end.
 In practice, we directly optimize a single global frequency bank $\textbf{W}_{1}$, rather than having the hypernetwork predict a separate $\textbf{W}_{1}$ for each low-resolution pixel.
-Not only does this better fit the idea of to represent the signal with a single, consistent basis; it also reduces the total parameter count and empirically leads to better results.
-Figure 4: Overview of Thera. A hypernetwork estimates the parameters of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
+Not only does this better fit the idea to represent the signal with a single, consistent basis, it also reduces the total parameter count.
 The described scheme, which we call Thera, is depicted in Fig.
-4.
-It allows for arbitrary-scale super-resolution, combining the proposed multi-scale signal representation with the expressivity of proven feature extraction backbones for SR and image restoration.
+3.
+It allows for arbitrary-scale super-resolution, combining the multi-scale signal representation within neural heat fields with the expressivity of proven feature extraction backbones for SR and image restoration.
 As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data.
-E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$.
+E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $t>\frac{1}{16}$.
 By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators.
@@
 Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.
-Prior at $\bm{t=0}$.
-To perform ASSR for out-of-domain scaling factors, we can use an unsupervised prior at $t=0$.
+3.3 Total Variation at $\bm{t=0}$
+To allow Thera to better generalize to higher, out-of-domain scaling factors, we can place an unsupervised regularizer at $t=0$.
 Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods.
-In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well.
-We use an $\ell^{1}$ variant of TV:
+In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well [11].
+We use an $\ell^{1}$ variant of TV,
 $\mathcal{L}_{\text{TV}}(\Phi(\mathbf{x},0))=\mathbb{E}_{\mathbf{x}}[|\nabla\Phi(\mathbf{x},0)|].$ (5)
-Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than falling back to a neighborhood approximation as in most previous work [38].
-Implementation and training.
-We have implemented Thera in JAX [7] to maximize the efficiency of the hypernetwork operations.
-A PyTorch [36] implementation to fit a Thera field to a single image is also provided.
-Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$.
+Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work [38].
+We further motivate this approach in Fig.
+5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
+3.4 Implementation and Training
+Thera is implemented in JAX [6].
+Similar to prior work [10, 24, 7, 56], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$.
 As corresponding targets, $48^{2}$ random pixels are sampled from the target patch.
-We train with the Adam optimizer [21] with batch size $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$.
+We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer [23] with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$.
 The learning rate is decayed to zero according to a cosine annealing schedule [30].
-As reconstruction loss, we minimize the mean absolute error (MAE), to which the TV loss from Eq.
+We use MAE as reconstruction loss, to which the TV loss from Eq.
 5 is added with a weight of $10^{-4}$.
-Training data is augmented by standard data augmentations, namely random flipping, rotation, and resizing.
-As in previous work [42, 24, 43], we implement geometric self-ensembling, where the results for four rotated versions of the input are averaged at test time.
-We did not observe a performance gain when also including reflections.
+Like previous work [43, 26, 44], we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF [10].
+In GSE, the results for four rotated versions of the input are averaged at test time.
+Including reflections did not improve performance.
 4 Results
-We evaluate Thera on a variety of datasets and with three popular feature extractors as backbones.
-We also conduct an ablation study that supports the design choices described in Sec.
-3.
-Throughout this section, we will evaluate three variants of our method which differ solely in the size of the hypernetwork and the number of field parameters:
-  • Thera-S: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and with only a single layer in hypernetwork, which maps features to field parameters.
+Throughout this section we evaluate three variants of our method, which differ solely in the size of the hypernetwork and the number of field parameters:
+  • Thera Air: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and the hypernetwork being a single $1\times 1$ convolution that maps features to field parameters.
 This version adds only 8,256 parameters on top of the backbone.
-  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork with 5 ConvNeXt layers.
-Its parameter count of 0.48 M approximately matches that of medium-sized competitors (MetaSR [17], LIIF [11], LTE [22] and CUF [43]).
-  • Thera-L: A large version with 512 components and a deeper hypernetwork of 16 ConvNeXt layers, for a total of 1.41 M additional parameters, comparable to CiaoSR [8].
+  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based [29] hypernetwork. Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like [7].
+  • Thera Pro: The strongest version uses a high-capacity, transformer-based [28, 25] hypernetwork.
+Its added parameter count is $\approx$4.63 M, still less than the most recent competitor [56] and much smaller than [9].
 Datasets and metrics.
 Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes.
-We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [51], BSDS100 [28] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [29] (high-resolution Japanese manga illustrations).
-Like prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets.
-Additional metrics are given in Appendix 0.C.
-Not all numbers could be computed for competing methods, because some do not provide public code or checkpoints, see Appendix 0.E.
+We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 [5], Set14 [53], BSDS100 [31], Urban100 [19], and Manga109 [32].
+Following prior work, we use peak signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric and compute it in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets.
+Additional quantitative results are given in Appendix C.
+Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix F).
 Backbones.
-Following related work, we combine our method with three standard backbones for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [52] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
-Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the DIV2K validation set.
-The highest PSNR value per backbone and scaling factor is in bold and the second highest is underlined.
-Backbone (params.) Method Num. of In-distribution Out-of-distribution
-add. params. $\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
-Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
-EDSR- baseline [24] (1.22 M) + interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
-MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
-LIIF [11] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
-LTE [22] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
-CUF [43] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
-CiaoSR [8] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
-CLIT [10] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
-Thera-S .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
-Thera-M 0.48 M 34.86 31.21 29.22 26.94 23.87 22.32 21.31 20.60
-Thera-L 1.41 M 34.88 31.22 29.24 26.96 23.89 22.33 21.32 20.61
-RDN [52] (22.0 M) + interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
-MetaSR [17] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
-LIIF [11] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
-LTE [22] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
-CUF [43] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
-CiaoSR [8] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
-CLIT [10] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
-Thera-S .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
-Thera-M 0.48 M 34.97 31.39 29.43 27.15 24.05 22.49 21.44 20.71
-Thera-L 1.41 M 34.98 31.40 29.45 27.16 24.06 22.49 21.45 20.72
-SwinIR [23] (12.0 M) + interpol. 35.09 31.42 29.48 27.17 24.08 22.53 21.52 20.82
-MetaSR [17] 0.45 M 35.15 31.40 29.33 26.94 23.80 22.26 21.26 20.54
-LIIF [11] 0.35 M 35.17 31.46 29.46 27.15 24.02 22.43 21.40 20.67
-LTE [22] 0.49 M 35.24 31.50 29.51 27.20 24.09 22.50 21.47 20.73
-CUF [43] 0.30 M 35.31 31.56 29.56 27.23 24.10 22.52
-CiaoSR [8] 1.43 M 35.28 31.53 29.56 27.26 24.12 22.52 21.48 20.59
-CLIT [10] 15.7 M 35.29 31.55 29.55 27.26 24.11 22.51 21.45 20.70
-Thera-S .008 M 35.27 31.57 29.58 27.25 24.13 22.55 21.52 20.77
-Thera-M 0.48 M 35.29 31.58 29.59 27.27 24.15 22.56 21.52 20.77
-Thera-L 1.41 M 35.29 31.59 29.61 27.30 24.17 22.59 21.54 20.79
+Like previous work, we combine our method with two standard backbones for super-resolution and image restoration: (i) EDSR-baseline [26] (1.22 M parameters) and (ii) RDN [54] (22.0 M parameters).
 4.1 Super-Resolution Performance
-We start by evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above.
-Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors.
-Thera achieves state-of-the-art performance.
-Thera-L outperforms all previous methods on 21 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline).
-Thera-M, which adds 3$\times$ fewer parameters to the backbone, still performs on par with or better than competitors that have comparable parameter counts.
-For scaling factors higher than ×$\times 2$, Thera-M also matches the performance of the more expensive, attention-based methods CiaoSR and CLIT, despite requiring 30×30$\times$30 × fewer parameters than CLIT.
-Notably, even our minimal version Thera-S, with the same anti-aliasing guarantees but with only 8,256 parameters on top of the backbone, reaches very competitive performance.
-It mostly outperforms all but the most parameter-intensive methods with the EDSR backbone.
-Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods, including attention-based ones, on 6 out of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (respectively 170×170$\times$170 × fewer than CiaoSR).
-Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors.
-We see this as rather strong evidence for the power of the principled multi-scale observation model that is baked into our network; especially for higher scaling factors.
-For a comparison with conventional interpolation, we also report numbers using Lanczos (sinc) resampling with the respective ×$\times 4$ backbone.
-This baseline is almost always outperformed by dedicated ASSR methods, indicating that the latter do learn scale-specific priors.
-Table 2: Evaluation on common benchmark datasets for in-distribution scale factors, with a SwinIR backbone.
-The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
-Method Set5 Set14 B100 Urban100 Manga109
-$\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4
-MetaSR 38.26 34.77 32.47 34.14 30.66 28.85 32.39 29.31 27.75 33.29 29.12 26.76 39.46 34.62 31.37
-LIIF 38.28 34.87 32.73 34.14 30.75 28.98 32.39 29.34 27.84 33.36 29.33 27.15 39.57 34.68 31.71
-LTE 38.33 34.89 32.81 34.25 30.80 29.06 32.44 29.39 27.86 33.50 29.41 27.24 39.63 34.79 31.79
-CUF 38.38 34.92 32.83 34.33 30.84 29.05 32.47 29.40 27.88 33.65 29.55 27.32
-CiaoSR 38.38 34.91 32.84 34.33 30.82 29.08 32.47 29.42 27.90 33.65 29.52 27.42 39.67 34.84 31.91
-CLIT 38.41 34.97 32.86 34.27 30.85 29.08 32.46 29.42 27.91 33.56 29.43 27.25
-Thera-S 38.32 34.88 32.75 34.30 30.85 29.07 32.46 29.39 27.88 33.61 29.56 27.25 39.62 34.93 31.91
-Thera-M 38.36 34.90 32.80 34.30 30.86 29.12 32.47 29.40 27.89 33.58 29.56 27.27 39.70 35.02 32.04
-Thera-L 38.41 34.93 32.82 34.33 30.87 29.11 32.49 29.42 27.91 33.67 29.62 27.33 39.71 35.01 32.07
-We further report the performance of Thera on five popular benchmark datasets in Tab.
-2.
-For these experiments, we use the SwinIR backbone.
-Overall, due to the unreasonably small data volume and out-of-distribution nature of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent.
-We posit that future work on ASSR would benefit from switching to a larger, statistically more meaningful test bed.
-Nonetheless, our methods perform best in 12 out of 15 cases.
-We once again observe that the performance of Thera-S is often comparable to that of methods with orders of magnitude higher parameter counts.
-When visually inspecting the super-resolution results – for examples see Fig.
-5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, particularly in the presence of repeating structures.
-This again highlights the advantage of hard-wiring a theoretically principled sampling model that rules out signal aliasing.
-For further qualitative results see Appendix 0.B.
+Low-res input LIIF [10] SRNO [45] MSIT [56] Thera Pro (ours) GT
+DIV2K
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
 DIV2K
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
-Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution.
-We use the SwinIR [23] backbone for all methods.
-Differences between methods are highlighted with bounding boxes.
-More examples are given in Appendix 0.B.
-4.2 Multi-Scale Consistency
-We are interested in learning a super-resolution model that is consistent across scales.
-For instance, if we super-resolve an image with a scaling factor $\times$4 then downsample it by a factor $\times$2, we expect the result to match what we get by directly super-resolving the original image by a factor $\times$2.
-We report in Tab.
-3 the quality of this match for several pairs of scaling factors, both for Thera with its in-built observation model and for three comparison methods where the observation model must be discovered from data.
-First, we notice that higher super-resolution quality does not correlate with better scale consistency, e.g., LTE in most cases achieves better consistency than CiaoSR.
-Second, while other methods generally exhibit good consistency, even for out-of-distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
-Table 3: Multi-scale consistency (in PSNR) of super-resolution methods across various pairs of scaling factors.
-$(2,4)$ $(2,6)$ $(2,12)$ $(2,24)$ $(3,12)$ $(3,24)$ $(4,24)$
-LIIF [11] 44.88 43.75 42.21 41.48 43.64 42.52 44.20
-LTE [22] 44.10 44.40 44.37 44.40 48.33 48.48 51.42
-CiaoSR [8] 44.24 43.31 42.10 41.56 41.47 41.03 45.47
-Thera-S 45.79 46.67 46.80 46.89 52.38 52.66 54.84
-Thera-M 45.26 46.10 46.34 46.41 51.91 52.16 54.64
-Thera-L 45.42 46.24 46.43 46.47 51.56 51.77 54.33
+Manga109
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN [54] backbone for all methods. Best viewed zoomed in.
+Qualitative results.
+Upon visual inspection – see Fig.
+4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures.
+Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Fig.
+9 in Appendix B).
+Backbone (params.) Method Num. of In-distribution Out-of-distribution
+add. params. $\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
+Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
+EDSR- baseline [26] (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
+MetaSR [18] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
+LIIF [10] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
+LTE [24] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
+CUF [44] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
+CiaoSR [7] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
+CLIT [9] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
+SRNO [45] 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
+MSIT [56] 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
+Thera Air (ours) .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
+Thera Plus (ours) 1.41 M 34.89 31.22 29.24 26.96 23.89 22.34 21.32 20.61
+Thera Pro (ours) 4.63 M 35.19 31.50 29.51 27.19 24.09 22.51 21.48 20.73
+RDN [54] (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
+MetaSR [18] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
+LIIF [10] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
+LTE [24] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
+CUF [44] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
+CiaoSR [7] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
+CLIT [9] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
+SRNO [45] 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
+MSIT [56] 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
+Thera Air .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
+Thera Plus 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
+Thera Pro 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
+Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the held-out DIV2K validation set.
+The highest PSNR value per backbone and scaling factor is bold and the second highest is underlined.
+Quantitative results.
+We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above.
+Table 1 shows PSNR values for all tested methods, for both in-distribution (×$\times 2$ to ×$\times 4$) and out-of-distribution (×$\times 6$ to ×$\times 30$) scaling factors.
+Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs.
+29.22 on EDSR ×$\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT [56], and less than a third of CLIT [9].
+Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count.
+This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction.
+For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective ×$\times 4$ backbone.
+This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
+Method Set5 Set14 B100 Urban100 Manga109
+$\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4 $\times$2 $\times$3 $\times$4
+MetaSR 38.22 34.63 32.38 33.98 30.54 28.78 32.33 29.26 27.71 32.92 28.82 26.55
+LIIF 38.17 34.68 32.50 33.97 30.53 28.80 32.32 29.26 27.74 32.87 28.82 26.68 39.26 34.21 31.20
+LTE 38.23 34.72 32.61 34.09 30.58 28.88 32.36 29.30 27.77 33.04 28.97 26.81 39.28 34.32 31.30
+CUF 38.28 34.80 32.63 34.08 30.65 28.92 32.39 29.33 27.80 33.16 29.05 26.87
+CiaoSR 38.29 34.85 32.66 34.22 30.65 28.93 32.41 29.34 27.83 33.30 29.17 27.11 39.51 34.57 31.57
+CLIT 38.26 34.80 32.69 34.21 30.66 28.98 32.39 29.34 27.82 33.13 29.04 26.91
+SRNO 38.32 34.84 32.69 34.27 30.71 28.97 32.43 29.37 27.83 33.33 29.14 26.98 39.52 34.67 31.61
+MSIT 38.31 34.85 32.72 34.26 30.70 28.97 32.42 29.35 27.81 33.27 29.14 26.93 39.44 34.62 31.58
+Thera Air 38.18 34.75 32.60 34.13 30.70 28.95 32.33 29.29 27.80 33.16 29.14 26.91 39.03 34.57 31.66
+Thera Plus 38.11 34.67 32.56 34.20 30.67 28.97 32.26 29.28 27.81 33.14 29.15 26.97 38.69 34.42 31.65
+Thera Pro 38.36 34.88 32.79 34.43 30.85 29.08 32.46 29.39 27.87 33.63 29.58 27.26 39.62 34.98 31.98
+Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN [54] backbone.
+The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
+Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Tab.
+2.
+Our method again outperforms all competing methods in all settings, often substantially (e.g., 29.58 vs.
+29.14 on Urban100 ×$\times 3$).
+We hypothesize that Thera’s hard-wired PSF is also beneficial when generalizing to unseen datasets.
+Once again we observe that the performance of Thera Air is often comparable to that of methods with orders of magnitude higher parameter counts.
+Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor [10].
+Due to its natural, Fourier-inspired representation, Thera is also infinitely differentiable, while ReLU-based competitors approximate the signal as a piecewise-linear function with null higher derivatives (last row).
+Fidelity of the Signal and its Derivatives.
+Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives [41].
+We observe similar effects for Thera, whose thermal activations at $t=0$ can be seen as a special case of periodic activations, cf.
+Figure 5.
+In fact, due to the use of thermal activations – and unlike all prior work based on multi-layer ReLU-activated fields – Thera is infinitely differentiable.
 4.2 Ablation Studies
-In Tab.
-4, we ablate individual components and design choices of our method to understand their contributions to overall performance.
-These comparisons use Thera-M with EDSR-baseline backbone, but are also representative of other settings.
+Experiment In-distribution Out-of-distribution
+$\times$2 $\times$3 $\times$4 $\times$6 $\times$18 $\times$30
+Thera Plus 34.89 31.22 29.24 26.96 22.34 20.61
+$\times$2 only, fixed $\kappa$ 35.01 31.00 28.84 26.47 21.97 20.33
+$\times$3 only, fixed $\kappa$ 34.74 31.25 29.25 26.91 22.25 20.54
+$\times$4 only, fixed $\kappa$ 34.47 31.17 29.24 26.97 22.32 20.59
+Fixed $\kappa$ 34.86 31.22 29.24 26.95 22.31 20.58
+No GSE 34.81 31.15 29.18 26.91 22.29 20.57
+No TV prior 34.89 31.23 29.24 26.87 20.42 18.68
+ReLU instead $\xi$ 34.80 31.11 29.13 26.87 22.28 20.57
+Predicted comps. 34.90 31.23 29.25 26.97 22.34 20.61
+Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
+In Table 3, we ablate individual components and design choices of our method to understand their contributions to overall performance.
+The comparisons use Thera Plus with EDSR backbone, and are representative of all variants.
 Single scale training.
-We ran three experiments using a single scale ($\times$2, $\times$3, or $\times$4) to test how this affects scale generalization.
-$\kappa$ was fixed for these experiments as multi-scale training is required to optimize it.
-As expected, we observe equal or even superior performance compared of single-scale training when tested at the training scale (marked in yellow in Tab.
-4), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
+We run three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization.
+$\kappa$ was fixed at the theoretically derived value for these experiments, as multi-scale training is required to optimize it.
+As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Tab.
+3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
 Trainable $\bm{\kappa}$.
 Fixing $\kappa$ at the theoretically derived value (Eq.
-5) leads to a small drop in performance.
-This suggests that there remain effects that are not accounted for by our proposed observation model, but that these are very minor.
+3) leads to a small drop in performance.
+This suggests that there remain effects that are not accounted for by our proposed observation model, albeit very minor.
 Geometric self-ensemble.
-In line with previous work [42, 24, 43] we see a notable performance boost with geometric self-ensembling (GSE).
+In line with previous work [43, 26, 44] we see a notable performance boost with geometric self-ensembling.
 Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
-Total variation. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
-No thermal activations.
+Total variation prior. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
+Thermal activations.
 We replace thermal activations (Eq.
-4) with standard ReLU activations.
+2) with standard ReLU activations.
 What remains is only the hypernetwork controlling the parameters of the local fields.
@@
 Shared components.
-We observe a noticeable performance drop, at all scales, if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$.
-Despite the greater flexibility afforded by individual per-field components, it seems challenging to learn them jointly with other, correlated parameters.
-Table 4: Ablation experiments with Thera-M using the EDSR-baseline backbone.
-Method In-distribution Out-of-distribution
-$\times$2 $\times$3 $\times$4 $\times$6 $\times$18 $\times$30
-Thera-M 34.86 31.21 29.22 26.94 22.33 20.60
-$\times$2 only, fixed $\kappa$ 34.98 30.96 28.81 26.40 21.93 20.30
-$\times$3 only, fixed $\kappa$ 34.71 31.22 29.22 26.88 22.24 20.53
-$\times$4 only, fixed $\kappa$ 34.44 31.14 29.22 26.95 22.30 20.58
-Fixed $\kappa$ 34.85 31.20 29.22 26.93 22.30 20.58
-No GSE 34.79 31.14 29.16 26.88 22.28 20.56
-No TV prior 34.87 31.21 29.21 26.86 20.82 19.12
-ReLU instead of $\xi$ 34.80 31.12 29.13 26.87 22.28 20.58
-Non-shared comps. 34.80 31.13 29.14 26.87 22.28 20.57
+Predicting $\bm{W}_{1}$ along with $\bm{b}_{1}$ and $\bm{W}_{2}$ leads to negligible gains.
+This comes at the cost of doubling the amount of field parameters that the hypernetwork predicts.
+Thus, Thera uses a shared, global frequency bank.
 4.3 Limitations and Future Work
-Thermal activations as introduced in this paper, and by extension our proposed ASSR method, come with relatively strict architectural requirements that only allow for a single hidden layer in the neural field.
-While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder.
+Neural heat fields as introduced in this paper, and by extension Thera, come with relatively strict architectural requirements that currently only allow for a single hidden layer in the neural field.
+While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder than necessary.
+Nonetheless, as our experiments show, the current neural heat field architecture does easily have enough capacity to model local, subpixel information for the scaling factor range discussed in this paper.
 We have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, and we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work.
 This could result in even higher parameter efficiency, and potentially better generalization.
-It may also be possible to construct similar formulations that do not require the usage of a hypernetwork, which could enable more streamlined training.
-Relative to other methods, Thera’s performance slightly drops for $\times$2 super-resolution when compared to higher scaling factors.
-We speculate that at small scaling factors Thera’s isotropic, continuous field representation is increasingly challenged by anisotropic effects from the oriented, discrete pixel grid.
-Further explorations are necessary to better understand this transition regime and perhaps come up with hybrid representations (which could potentially benefit also other methods at the interface of gridded and implicit representations like Instant-NGP [33]).
-Finally, we expect that more advanced priors than TV could be even more effective at regularizing $\Phi$.
-Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.
+We also expect that more advanced priors than TV could be even more effective at regularizing $\Phi$.
+Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR quality for all scaling factors.
 5 Conclusion
 We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations.
-Our method, Thera, implicitly describes an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) analytically, with negligible overhead.
-Our experimental evaluation shows that Thera outperforms competing methods in most cases, while at the same time being more parameter efficient, and offering theoretical guarantees.
-We are convinced that Thera-style representations could also benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful observation models like our Gaussian PSFs.
-Appendix 0.A Theory
-0.A.1 Preliminaries
-As was described in Sec.
-3.1, the idea underlying our neural field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
+Our proposed neural heat fields implicitly describe an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) between scales analytically, with negligible overhead.
+Our experimental evaluation shows that Thera, our ASR method based on neural heat fields, consistently outperforms competing methods.
+At the same time, it is more parameter-efficient and offers theoretical guarantees w.r.t.
+aliasing.
+We believe that Thera-style representations could benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful and theoretically grounded observation models.
+Appendix A Theory
+A.1 Preliminaries
+As was described in Section 3.1, the idea underlying our neural heat field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
 $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;.$ (6)
@@
 For all the formulations here, we define the image $I$ to correspond to the coordinates $x_{1},x_{2}\in[-0.5,0.5]$.
-0.A.2 Thermal Diffusivity Coefficient
+A.2 Thermal Diffusivity Coefficient
 To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$.
@@
 $\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}N^{2}}.$ (12)
-For the Thera formulation presented in Sec.
-3.2, we want $\Phi$ to contain a single pixel at $t=1$.
+For our proposed Thera formulation, we want $\Phi$ to contain a single pixel at $t=1$.
 This is the pixel from the low-resolution input which will become $SR^{2}$ pixels for super-resolution with a scaling factor of $SR$.
@@
 Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the formulas for a Gaussian filter with the appropriate number of dimensions.
-0.A.3 Relationship Between $t$ and $s$
+A.3 Relationship Between t and s
 Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate.
@@
 10 and Eq.
-13, we can easily find the quadratic relationship
+11, we can easily find the quadratic relationship
 $t=S^{2}.$ (14)
-For instance, if we want to upsample the image by a factor of 2, we should use
-$t=0.5^{2}=0.25.$ (17)
+For instance, if we want to upsample the image by a factor of 2, we should use $t=0.5^{2}=0.25$.
 Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling.
 This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
-0.A.4 Other Filters
-In theory, the formulation presented in Sec.
-3.1 allows us to use any low-pass filter we want, since we can modulate different components freely.
-Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation.
+Image domain
+Refer to caption Refer to caption Refer to caption Refer to caption
+Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA
+Neural field
+Refer to caption Refer to caption Refer to caption Refer to caption
+$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$
+Figure 6: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with neural heat fields (bottom row), which is computed without convolutions or over-sampling.
+The heat field was supervised at $\Phi(\mathbf{x},1)$ only.
+Figure 6 shows an example where we fit a neural heat field at $t=1$ to the image.
+After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq.
+4.
+We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
+$t=0$
+Refer to caption Refer to caption Refer to caption Refer to caption
+$t=1$
+Refer to caption Refer to caption Refer to caption Refer to caption
+a) b) c) d)
+Figure 7: Example situations where aliasing would occur without the suppression of high frequencies as modeled by neural heat fields.
+Sampling the center location at $t=0$ would not be representative of the pixel’s footprint.
+Figure 8: Blurring an image before a point-wise sample is equivalent to observing with PSF equivalent to the blur kernel.
+In Fig.
+7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations.
+Note that aliasing is not always as obvious to the eye as Moiré patterns: in the cases shown in the figure, it would simply mean that sampling the center location at $t=0$ would not be representative of the pixel’s footprint.
+Figure 8 further illustrates how such blur is equivalent to a scale-appropriate anti-aliasing filter.
+A.4 Other Filters
+In theory, the formulation presented in Section 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely.
+Gaussian filters are an obvious choice since they are often used as anti-aliasing filters, and since they are fully defined by a single parameter, the standard deviation.
 Initial explorations of a sharp low-pass filter that completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely due to the associated effect on gradients during training.
-It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way.
+It remains an open question whether more complex filters (e.g., Butterworth) would improve the current formulation in any way.
 For quantitative evaluations, this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios, this may be desirable.
-0.A.5 Initialization of Components
+A.5 Initialization of Components
 We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Eq.
-2, is important.
+1, is important.
 Sitzmann et al.
@@
 up to a given maximum frequency, allotting more components to higher frequencies. See code for more details.
-Appendix 0.B Additional Qualitative Results
-Figures 9 and 10 contain additional qualitative results for lower and higher scaling factors, respectively.
-Again, we see that Thera can reconstruct the targets more faithfully.
-This is particularly (but not exclusively) visible in repetitive structural elements.
-It seems evident that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this.
-Furthermore, we provide a showcase of the continuous upsampling capabilities of our method in Fig.
-6.
-Figure 6: Showcase of multiscale upsampling using Thera-L with a SwinIR [23] backbone, shown with non-integer scaling factors.
-Appendix 0.C Additional Quantitative Results
-Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab.
-1.
-We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric.
-Furthermore, Tab.
-6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
-Finally, in Fig.
-7 we compare the number of additional parameters and PSNR values for various methods and upsampling factors on DIV2K.
-Figure 7: Plots depicting the number of additional parameters vs.
-PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 2$ (left), ×$\times 3$ (center) and ×$\times 4$ (right) upsampling tasks.
-Table 5: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the SwinIR [23] backbone for all models.
-In-distribution Out-of-distribution
-×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$$\times 12$$\times 18$$\times 24$$\times 30$
-LIIF [11] .942 .883 .830 .753 .702 .641 .594 .568 .551
-LTE [22] .942 .883 .832 .755 .704 .643 .595 .569 .552
-CiaoSR [8] .942 .883 .832 .756 .706 .645 .598 .572 .551
-Thera-S .943 .884 .833 .757 .706 .644 .596 .569 .552
-Thera-M .943 .885 .834 .758 .707 .645 .596 .569 .552
-Thera-L .942 .885 .834 .759 .708 .646 .597 .570 .552
-Table 6: Evaluation on common benchmark datasets for out-of-distribution scale factors, using the SwinIR backbone.
-The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
+Appendix B Continuous Upsampling Example
+In Fig. 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
+Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN [54] backbone, shown with non-integer scaling factors.
+Appendix C Additional Quantitative Results
+C.1 Further Metrics
+In-distribution OOD
+×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$
+LIIF [10] .934 .866 .804 .704 .636
+LTE [24] .936 .868 .807 .707 .639
+CiaoSR [7] .938 .871 .814 .718 .651
+SRNO [45] .937 .873 .819 .738 .685
+MSIT [56] .942 .882 .829 .751 .700
+Thera Pro .943 .884 .833 .757 .706
+Table 4: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set.
+We use the RDN backbone for all models.
+Some methods did not provide checkpoints at the time of writing, see Appendix F.
 Method Set5 Set14 B100 Urban100 Manga109
-$\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12
-MetaSR 29.09 27.02 24.82 26.58 25.09 23.33 25.94 24.87 23.59 24.16 22.75 21.31 27.29 24.96 22.35
-LIIF 29.46 27.36 26.82 25.34 26.07 25.01 24.59 23.14 27.69 25.28
-LTE 29.50 27.35 26.86 25.42 26.09 25.03 24.62 23.17 27.83 25.42
-CUF 29.57 26.91 26.11 24.69
-CiaoSR 29.62 27.45 24.96 26.88 25.42 23.38 26.13 25.07 23.68 24.84 23.34 21.60 28.01 25.61 22.79
-CLIT 29.69 27.62 26.94 25.55 26.15 25.09 24.77 23.33
-Thera-S (ours) 29.45 27.33 25.00 26.85 25.40 23.40 26.11 25.05 23.69 24.68 23.23 21.55 27.94 25.49 22.73
-Thera-M (ours) 29.47 27.43 25.07 26.90 25.41 23.40 26.13 25.07 23.69 24.70 23.25 21.57 28.05 25.59 22.79
-Thera-L (ours) 29.48 27.46 25.09 26.94 25.45 23.42 26.14 25.08 23.71 24.79 23.32 21.62 28.13 25.69 22.85
-Appendix 0.D Hypernetwork Architecture
-In Fig.
-8, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L.
-The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image.
-It is then followed by a number of (residual) ConvNeXt [26] blocks.
-Each block consists of (i) a $3\times 3$ convolution that maps to $d$-dimensional features, (ii) a layer normalization [2] operation, (iii) a $1\times 1$ convolution with $4d$ filters, (iv) a GELU [15] non-linearity, and (v) another $1\times 1$ convolution with $d$ filters.
-Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al.
-[26] – to $3\times 3$ to reduce the number of parameters.
-Furthermore, note that in Thera-L, the first convolution per block uses $d$ groups (effectively implementing a channel-wise convolution), while for the two smaller variants we used full convolutions, i.e., a single group.
-In our implementation, Thera-S uses zero ConvNeXt blocks, Thera-M uses 5 blocks with $d=64$, and Thera-L uses 6 blocks with $d=64$ followed by 7 blocks with $d=96$ and 3 blocks with $d=128$.
-For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
-Figure 8: Schematics of the hypernetworks for a) Thera-S, b) Thera-M and c) Thera-L.
-Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [40].
-There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L.
-Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork.
-The total parameter count of the hypernetwork is 8,192, 0.48 M, and 1.41 M for the three variants, respectively.
+$\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8
+MetaSR 29.04 29.96 26.51 24.97 25.90 24.83 23.99 22.59
+LIIF 29.15 27.14 26.64 25.15 25.98 24.91 24.20 22.79 27.33 25.04
+LTE 29.32 27.26 26.71 25.16 26.01 24.95 24.28 22.88 27.49 25.12
+CUF 29.27 26.74 26.03 24.32
+CiaoSR 29.46 27.36 26.79 25.28 26.07 25.00 24.58 23.13 27.70 25.40
+CLIT 29.39 27.34 26.83 25.35 26.07 25.00 24.43 23.03
+SRNO 29.38 27.28 26.76 25.26 26.04 24.99 24.43 23.02 27.66 25.31
+MSIT 29.34 27.29 26.75 25.26 26.05 24.98 24.43 22.99 27.61 25.26
+Thera Air 29.31 27.25 26.76 25.27 26.05 24.99 24.39 23.00 27.70 25.30
+Thera Plus 29.31 27.29 26.80 25.32 26.07 25.00 24.45 23.04 27.80 25.41
+Thera Pro 29.51 27.34 26.90 25.38 26.12 25.04 24.70 23.24 27.94 25.49
+Table 5: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN [54] backbone.
+For some methods, code and/or checkpoints were not publicly avalibale, see Appendix F.
+Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1.
+We use the SSIM implementation from torchmetrics [13].
+We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
+Furthermore, in Table 5 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
+C.2 Parameter Efficiency
+In Fig. 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
+$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$34.734.734.734.734.834.834.834.834.934.934.934.93535353535.135.135.135.135.235.235.235.2Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×2 Scaling Factor
+$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$3131313131.131.131.131.131.231.231.231.231.331.331.331.331.431.431.431.431.531.531.531.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×3 Scaling Factor
+$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$2929292929.129.129.129.129.229.229.229.229.329.329.329.329.429.429.429.429.529.529.529.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×4 Scaling Factor
+Figure 10: Comparison of ASR methods for different scaling factors (×2, ×3, and ×4) on DIV2K. Thera consistently achieves better performance at lower parameter counts across all scaling factors.
+Appendix D Hypernetwork Architecture
+In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN [41].
+Thera Plus uses 6 ConvNeXt [29] blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer.
+Projection blocks are added between blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
+For Thera Pro, two SwinIR [25] blocks are used with depth 7 and 6, respectively, and 6 attention heads per layer.
+There are 128 field parameters for Thera Air and 2048 for the larger variants Thera Plus and Thera Pro.
+These numbers are computed as follows: Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork.
+The total parameter count of the hypernetwork is 8,192, 1.41 M, and 4.63 M for the three variants, respectively.
 To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters.
 We highlight that these do not come from the hypernetwork.
-Appendix 0.E Replication of Other Methods
-We encountered challenges attempting to recreate results of some of the competing methods, which explains why some numbers are missing in various tables or differ from the originally reported ones.
+Appendix E Real-World Optical Zoom Data
+To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset [14].
+We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) [21] that estimates just 6 parameters per image to correct domain-specific affine distortions.
+Despite its name, the STN is not a transformer network in the modern sense that employs attention mechanisms; rather, it is an image model block that explicitly allows the spatial manipulation of data within a convolutional neural network.
+The STN component follows a standard architecture, consisting of a simple convolutional localization network with adaptive pooling to handle variable input sizes.
+The network processes the input image along with the scale factor and outputs six parameters of a 2D affine transformation ($xy$-translation, anisotropic $xy$-scale, rotation and shear).
+With approximately 10K parameters, this lightweight component efficiently corrects for geometric distortions while adding minimal computational overhead.
+Thera++ addresses an important limitation of the COZ dataset.
+When looking at the data it becomes obvious that COZ samples are not perfectly aligned, resulting in $xy$-jitter between images.
+Additionally, dynamic objects like people, dust, leaves, and moving shadows appear inconsistently across images of the same scene, significantly increasing the noise level.
+These challenges make super-resolution particularly difficult for this dataset.
+However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Tab.
+6.
+Method In-distribution Out-of-distribution
+$\times$2 $\times$3 $\times$4 $\times$5 $\times$6
+MetaSR [18] 28.70 26.55 25.17 24.31 23.25
+LIIF [10] 28.72 26.61 25.16 24.32 23.23
+LTE [24] 28.67 26.55 25.15 24.37 23.26
+SRNO [45] 28.73 26.59 25.15 24.31 23.25
+LIT [9] 28.74 26.58 25.15 24.35 23.19
+LMI [14] 28.86 26.66 25.22 24.39 23.29
+Thera++ 29.06 26.84 25.45 24.49 23.46
+Table 6: Results (PSNR in dB) on the COZ test set.
+Appendix F Reproducibility of Existing Methods
+We encountered challenges attempting to recreate the results reported for some of the competing methods, which explains why some are missing or differ from the originally reported numbers.
 Details are provided below.
@@
 We have denoted those missing values with “—” in the tables.
-Furthermore, we could not create any qualitative samples and could not include CUF in the multi-scale consistency study in Sec.
-4.2.
+Furthermore, we could not create any qualitative samples using CUF.
 CLIT [9].
-For CLIT, at the time of writing, there is public code available, however, checkpoints are not.
-We attempted to recreate their models, but due to the cascaded training schedule and the big model size, the training process would require excessive amounts of GPU time: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090s.
-The authors did unfortunately not respond to our requests for the trained checkpoints used in their paper.
+For CLIT, at the time of writing, there is a public code, but no checkpoints.
+We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090 GPUs.
+Unfortunately, the authors did not respond to our requests for the trained checkpoints used in their paper.
 CiaoSR [7].
-We noticed that in the official CiaoSR implementation, border cropping prior to evaluation on the DIV2K dataset deviated slightly from all other methods.
-We aligned this to enable a meaningful comparison and re-computed all DIV2K numbers.
-This explains the slight deviation from the original paper.
-DIV2K ×$\times 2$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-DIV2K ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-DIV2K ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 2$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109 ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109 ×$\times 3$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
-Figure 9: Further qualitative examples for ×$\times 2$ and ×$\times 3$ upsampling factors. Best viewed zoomed in.
-DIV2K ×$\times 4$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-DIV2K ×$\times 4$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 4$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 4$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100 ×$\times 6$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109 ×$\times 6$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109 ×$\times 6$
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
-Figure 10: Further qualitative examples for ×$\times 4$ and ×$\times 6$ upsampling factors. Best viewed zoomed in.
-Appendix 0.F Computational Requirements
-For comparison of computational requirements between methods, we provide training time, inference time, and inference VRAM requirements in Tab.
-7, for the EDSR-baseline backbone.
-These values assume NVIDIA GeForce RTX 3090 Ti GPUs, and they show that we are more efficient than recent transformer-based methods, but still reach superior performance, as shown in the main paper.
-Table 7: Computational requirements of various methods.
-w/ EDSR-b Training time Inference time VRAM
-(GPU days)
-MetaSR 0.42 3.68 ms 588 MiB
-LIIF 0.71 13.7 ms 1174 MiB
-LTE 0.79 17.1 ms 1250 MiB
-CiaoSR $\sim 2.6$ 37.7 ms 6474 MiB
-CLIT $\sim 4.6$ 16.0 ms 2984 MiB
-Thera-S 1.0 3.52 ms 754 MiB
-Thera-M 2.2 10.0 ms 2290 MiB
-Thera-L 3.3 14.3 ms 2292 MiB
+We found that in the official CiaoSR implementation border cropping prior to evaluating on DIV2K deviated slightly from all other methods.
+We thus adapted the evaluation code to match the competition and enable a meaningful comparison.
+All DIV2K numbers were re-computed with the corrected code, resulting in slight deviations from those reported in the original paper.
+MSIT [56].
+The numbers reported in the paper were achieved by training on a roughly 3$\times$ larger training set, comprising not only DIV2K but also Flickr2K.
+For a fair comparison with all other ASR methods used in our paper, we re-trained MSIT on DIV2K alone.
+We used the authors’ official code and configuration files and trained for two stages, each comprising 1050 epochs.
+For the second (RiM) training stage we used scaling factors between ×$\times 1$ and ×$\times 4$.
+Performance generally drops as a result of the smaller training set, in line with the ablation experiments reported for MSIT, where the models were also trained only with DIV2K.

### v4

@@
 The project page is at https://therasr.github.io.
+1 Introduction
+Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results.
+However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor.
+Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility (Hu et al., 2019).
+Notably, with LIIF, Chen et al.
+(2021) pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors.
+LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields (Lee & Jin, 2022; Cao et al., 2023; Chen et al., 2023; Zhu et al., 2025).
+This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics (Xie et al., 2022).
+By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
 Figure 1: We present Thera, the first method for arbitrary-scale super-resolution with a built-in physical observation model.
@@
 The field’s architecture automatically attenuates frequencies as a function of the scaling factor so as to match the output resolution at which the signal is re-sampled.
-1 Introduction
-Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results.
-However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor.
-Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility [18].
-Notably, with LIIF, Chen et al.
-[10] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors.
-LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields [24, 7, 9, 56].
-This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [48].
-By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
 While neural fields naturally model continuous functions, they do not easily allow for observations of such functions other than point-wise evaluations.
@@
 This explains the initially counterintuitive relevance of anti-aliasing for super-resolution: When using neural fields, signals are first upsampled to infinite (continuous) resolution and then resampled at the desired resolution, and this latter operation must be done carefully.
-Incorporating a physically plausible observation model is not trivial [2, 3, 27, 51, 17, 4], but has the potential to avoid aliasing.
+Incorporating a physically plausible observation model is not trivial (Barron et al., 2021; 2022; Lindell et al., 2022; Yang et al., 2022; Hu et al., 2023; Barron et al., 2023), but has the potential to avoid aliasing.
 For this reason, Chen et al.
-[10] and successor works [24, 7, 9, 56] have already taken a first step towards learning multi-scale representations, via cell encoding.
-Fundamentally, these “learning-based anti-aliasing” approaches try to learn an integrated version of the field for any scaling factor, wasting field capacity for approximating a relation that can be described exactly through Fourier theory.
-$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$31.531.531.531.531.631.631.631.631.731.731.731.731.831.831.831.831.931.931.931.93232323232.132.132.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
-Figure 2: Comparison of recent ASR methods, averaged over ×$\times\{2,3,4\}$ scales.
+(2021) and successor works (Lee & Jin, 2022; Cao et al., 2023; Chen et al., 2023; Zhu et al., 2025) have already taken a first step towards learning multi-scale representations, via cell encoding.
+Fundamentally, these “learning-based anti-aliasing” approaches require the scaling factor (or, equivalently, the output pixel area) as additional input to the neural field and learn an integrated (i.e., appropriately blurred and therefore anti-aliased) version of the field for each scaling factor; arguably wasting field capacity to approximate a relation that can be described exactly through Fourier theory.
+$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$31.531.531.631.631.731.731.831.831.931.9323232.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
+Figure 2: Comparison of recent ASR methods, averaged over $\times\{2,3,4\}$ scales.
 We generally achieve higher performance at lower parameter counts.
@@
 In this work, we combine recent advances in implicit neural representations with ideas from classical signal theory to introduce neural heat fields, a novel type of neural field that guarantees anti-aliasing by construction.
-The key insight is that sinusoidal activation functions [41] enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory.
+The key insight is that sinusoidal activation functions (Sitzmann et al., 2020b) enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory.
 This allows for the exact computation of Gaussian-blurred versions of the field for any desired (isotropic) blur radius.
 When rasterizing an image, the field can therefore be queried with a Gaussian PSF that matches the target resolution, effectively preventing aliasing.
+In practice, heat fields receive an additional input coordinate $t$, controlling the strength of the Gaussian blur applied to the signal.
+Unlike learning-based anti-aliasing, the resulting filtering operation is expressed analytically, rather than learned from data.
+In other words, previous approaches fit a 3D field ($x$, $y$, and scalescale) while we only need to fit a 2D field ($x$ and $y$), whereas the scale dimension is computed analytically, significantly reducing field complexity and data requirements.
 Notably, filtering with neural heat fields incurs no computational overhead: The querying cost is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
-Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork [16] with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Fig.
-1).
-Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Fig.
-2).
+Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork (Ha et al., 2017) with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Figure 1).
+Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Figure 2).
 To the best of our knowledge, Thera is also the first neural field method to allow bandwidth control at test time.
@@
 A neural field, also called an implicit neural representation, is a neural network trained to map coordinates onto values of some physical quantity.
-Recently, neural fields have been used for parameterizing various types of visual data, including images [22, 41, 42, 10, 24, 12, 47], 3D scenes (e.g., represented as signed distance fields [36, 41, 40, 46, 47], occupancy fields [33, 37], LiDAR fields [20], view-dependent radiance fields [34, 2, 3, 4, 47]), or digital humans [52, 55, 8, 49, 15].
+Recently, neural fields have been used for parameterizing various types of visual data, including images (Karras et al., 2021; Sitzmann et al., 2020b; Tancik et al., 2020; Chen et al., 2021; Lee & Jin, 2022; de Lutio et al., 2019; Wu et al., 2023), 3D scenes (e.g., represented as signed distance fields (Park et al., 2019; Sitzmann et al., 2020b; a; Williams et al., 2022; Wu et al., 2023), occupancy fields (Mescheder et al., 2019; Peng et al., 2020), LiDAR fields (Huang et al., 2023), view-dependent radiance fields (Mildenhall et al., 2021; Barron et al., 2021; 2022; 2023; Wu et al., 2023)), or digital humans (Yenamandra et al., 2021; Zheng et al., 2022; Cao et al., 2022; Xiu et al., 2022; Giebenhain et al., 2023).
 Frequently, it is desirable to impose some prior over the space of learnable implicit representations.
-A common approach for such conditioning is encoder-based inference [48], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [10, 24, 44, 7, 9].
-The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [16], mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$.
-An early example of this approach, which is gaining popularity [48], was proposed in [41].
+A common approach for such conditioning is encoder-based inference (Xie et al., 2022), where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local (Chen et al., 2021; Lee & Jin, 2022; Vasconcelos et al., 2023; Cao et al., 2023; Chen et al., 2023).
+The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork (Ha et al., 2017), mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$.
+An early example of this approach, which is gaining popularity (Xie et al., 2022), was proposed in Sitzmann et al.
+(2020b).
 2.2 Arbitrary-Scale Super-Resolution
 ASR is the sub-field of single-image SR in which the desired SR scaling factor can be chosen at inference time to be (theoretically) any positive number, allowing maximum flexibility, such as that of interpolation methods.
-The first work along this line is MetaSR [18], which infers the parameters of a convolutional upsampling layer using a hypernetwork [16] conditioned on the desired scaling factor.
-An influential successor work is LIIF [10], in which the high-resolution image is implicitly described by local neural fields.
+The first work along this line is MetaSR (Hu et al., 2019), which infers the parameters of a convolutional upsampling layer using a hypernetwork conditioned on the desired scaling factor.
+An influential successor work is LIIF (Chen et al., 2021), in which the high-resolution image is implicitly described by local neural fields.
 These fields are conditioned via concatenation, with features extracted from the low-resolution input image.
@@
 Most subsequent work has since been built upon the LIIF framework.
-For example, UltraSR [50] improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields [34, 2, 3, 4].
-LTE [24] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP.
+For example, UltraSR (Xu et al., 2021) improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields (Mildenhall et al., 2021; Barron et al., 2021; 2022; 2023).
+LTE (Lee & Jin, 2022) makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP.
 Vasconcelos et al.
-[44] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling.
-More recently, methods like CiaoSR [7], CLIT [9], and most recently MSIT [56] have integrated (multi-scale) attention mechanisms, improving reconstruction quality.
-In a parallel line of research, Wei et al.
-[45] propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
+(2023) use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling.
+More recently, methods like CiaoSR (Cao et al., 2023), CLIT (Chen et al., 2023), and most recently MSIT (Zhu et al., 2025) have integrated (multi-scale) attention mechanisms, improving reconstruction quality.
+In a parallel line of research, Wei & Zhang (2023) propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
+Another line of work employs generative models such as denoising diffusion for SR (Saharia et al., 2022; Gao et al., 2023).
+While most methods minimize per-pixel errors (essentially predicting the minimum mean square error estimate), generative models are trained to produce more realistic-looking outputs by predicting one of many plausible high-resolution images.
+However, since specific ground truth details are not exactly recovered, such models typically report worse distortion metrics (like PSNR and SSIM) compared to pixel-based methods, c.f.
+Blau & Michaeli (2018); Delbracio & Milanfar (2023).
+In this paper, we adopt a pixel-based objective to preserve fidelity to the ground truth, which is important for many downstream applications (e.g., face or license plate recognition).
 2.3 Anti-Aliasing in Neural Fields
@@
 Barron et al.
-[2] proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF [34].
-Improvements were later proposed for unbounded scenes [3] and to improve efficiency [17].
-In [4], Barron et al.
-tackle anti-aliasing within the Instant-NGP [35] approach.
-Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [27], polynomial neural fields [51] or cascaded training [39], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks.
+(2021) proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF (Mildenhall et al., 2021).
+Improvements were later proposed for unbounded scenes (Barron et al., 2022) and to improve efficiency (Hu et al., 2023).
+Barron et al.
+(2023) tackle anti-aliasing within the Instant-NGP (Müller et al., 2022) approach.
+Recent work has succeeded in limiting the bandwidth using multiplicative filter networks (Lindell et al., 2022), polynomial neural fields (Yang et al., 2022) or cascaded training (Shabanov et al., 2024), although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks.
 These methods are not a good fit for ASR because they do not allow for continuous anti-aliasing, nor bandwidth control at test time.
@@
 3 Method
+In this section we introduce Thera, a novel neural fields-based ASR method that guarantees analytical anti-aliasing at any desired output resolution at no additional cost.
+First, we present neural heat fields, a special type of neural field that inherently achieves anti-aliasing by implicitly attenuating high-frequency components as a function of a time coordinate.
+Next, we propose a mechanism for learning a prior over a grid of neural heat fields, enabling them to represent a multi-scale output image conditioned on a lower-resolution input image.
+Finally, we show that our formulation allows us to impose a regularizer on the underlying, continuous signal itself – something that, to the best of our knowledge is not possible in previous methods.
 Figure 3: Overview of Thera.
@@
 This field is then rasterized at the appropriate sampling rate (resolution) to yield a part of the final output image (red square).
-In this section we introduce Thera, a novel neural fields-based ASR method that guarantees analytical anti-aliasing at any desired output resolution at no additional cost.
-First, we present neural heat fields, a special type of neural field that inherently achieves anti-aliasing by implicitly attenuating high-frequency components as a function of a time coordinate.
-Next, we propose a mechanism for learning a prior over a grid of neural heat fields, enabling them to represent a multi-scale output image conditioned on a lower-resolution input image.
-Finally, we show that our formulation allows us to impose a regularizer on the underlying, continuous signal itself – something that, to the best of our knowledge is not possible in previous methods.
 3.1 Neural Heat Fields for Analytical Anti-Aliasing
@@
 One must therefore apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$.
-The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency and a time coordinate $t$.
-This behavior is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
+The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency as well as a time coordinate, which we call $t$.
+The time coordinate acts as a third, continuous input to the neural field and controls the amount of re-scaling, and therefore the Gaussian blur applied to the signal.
+This perfectly mimics how high-frequency components decay faster than low-frequency ones in the analytical solution to the heat equation.
+The detailed derivation can be found in Appendix A.
+The behavior described above is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
 $\Phi(\mathbf{x},t)$ $=\mathbf{W}_{2}\cdot\xi\left(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1},\nu(\mathbf{W}_{1}),\kappa,t\right)+\textbf{b}_{2},$ (1)
@@
 Here, $\left|\nu\right|=\left|\nu(\mathbf{W}_{1})\right|$ denotes the row-wise Euclidean norm of $\mathbf{W}_{1}$, representing the magnitudes of the implied wave numbers (frequencies).
-Interestingly, Eq.
-1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A.
+Interestingly, Equation 1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A.
 We therefore refer to this MLP as a neural heat field.
@@
 To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$.
-The relationship between the cut-off frequency of the filter and $t$ is controlled by the diffusivity constant $\kappa$, which we can freely set to any positive number.
-For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$.
+The relationship between the cut-off frequency of the filter and $t$ is controlled by a global diffusivity constant $\kappa$, which defines how fast components of different frequencies decay over time in the underlying PDE model.
+We can freely set $\kappa$ to any positive number, but for simplicity, and without loss of generality, we set $\kappa$ in our theoretical derivations such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ corresponds to $t=1$.
 Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
-$\kappa=\frac{\sqrt{\ln(4)}}{2f_{s}\pi^{2}}.$ (3)
+$\kappa=\frac{\ln(4)}{2f_{s}^{2}\pi^{2}}.$ (3)
 To subsample the signal $\mathcal{D}$ by a factor $S$, the field $\Phi$ should be sampled at
-$t=S^{2}.$ (4)
-For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Appendix A.
+$t=S^{2},$ (4)
+where $S$ is the subsampling rate, i.e., the inverse of the scaling factor.
+In other words, the equation above defines the correct value for $t$ according to the scaling factor at which the field should be sampled.
+For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Section A.
 3.2 Learning a Super-Resolution Prior
@@
 First, our formulation restricts the choice of architecture to MLPs with a single hidden layer.
-Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($10<t<$ 1) remains ill-posed.
+Second, while it theoretically guarantees the downsampling operation ($t>1$), the upsampling operation ($0<t<1$) remains ill-posed.
 To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data.
-Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, mapping the low-resolution input image to the $N$ parameters of each local neural heat field.
-As originally proposed by LIIF [10] and adopted by recent ASR methods [24, 7, 44, 9], each local field spans the area of one pixel of the low-resolution input.
+Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times C}\to\mathbb{R}^{W\times H\times N}$.
+First, a standard backbone, as used in previous work (Chen et al., 2021; Lee & Jin, 2022; Vasconcelos et al., 2023; Cao et al., 2023; Chen et al., 2023), extracts image features from the low-resolution input image.
+Then, the hypernetwork maps these features to the $N$ parameters of each local neural heat field.
+As originally proposed by LIIF (Chen et al., 2021) and adopted by recent ASR methods (Lee & Jin, 2022; Cao et al., 2023; Vasconcelos et al., 2023; Chen et al., 2023), each local field spans the area of one pixel of the low-resolution input.
 It is important that even though the fields themselves model only a local part of the image, the hypernetwork informs them with contextual features collected over a large receptive field.
-The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [10, 24, 44, 7, 9], followed by one or more layers that produce feature maps of depth $N$.
 During training, the local fields are supervised with values of high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$ (ensuring that the signal is correctly blurred for the target resolution), and the entire architecture is optimized end-to-end.
@@
 Not only does this better fit the idea to represent the signal with a single, consistent basis, it also reduces the total parameter count.
-The described scheme, which we call Thera, is depicted in Fig.
-3.
+The described scheme, which we call Thera, is depicted in Figure 3.
 It allows for arbitrary-scale super-resolution, combining the multi-scale signal representation within neural heat fields with the expressivity of proven feature extraction backbones for SR and image restoration.
@@
 Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods.
-In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well [11].
+In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well (Chugunov et al., 2024).
 We use an $\ell^{1}$ variant of TV,
 $\mathcal{L}_{\text{TV}}(\Phi(\mathbf{x},0))=\mathbb{E}_{\mathbf{x}}[|\nabla\Phi(\mathbf{x},0)|].$ (5)
-Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work [38].
-We further motivate this approach in Fig.
-5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
+Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work (Rudin et al., 1992).
+We further motivate this approach in Figure 5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
 3.4 Implementation and Training
-Thera is implemented in JAX [6].
-Similar to prior work [10, 24, 7, 56], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$.
+Thera is implemented in JAX (Bradbury et al., 2018).
+Similar to prior work (Chen et al., 2021; Lee & Jin, 2022; Cao et al., 2023; Zhu et al., 2025), we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$.
 As corresponding targets, $48^{2}$ random pixels are sampled from the target patch.
-We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer [23] with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$.
-The learning rate is decayed to zero according to a cosine annealing schedule [30].
+We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer (Kingma & Ba, 2015) with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$.
+The learning rate is decayed to zero according to a cosine annealing schedule (Loshchilov & Hutter, 2016).
 We use MAE as reconstruction loss, to which the TV loss from Eq.
 5 is added with a weight of $10^{-4}$.
-Like previous work [43, 26, 44], we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF [10].
+Like previous work (Timofte et al., 2016; Lim et al., 2017; Vasconcelos et al., 2023), we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF (Chen et al., 2021).
 In GSE, the results for four rotated versions of the input are averaged at test time.
@@
 This version adds only 8,256 parameters on top of the backbone.
-  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based [29] hypernetwork. Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like [7].
-  • Thera Pro: The strongest version uses a high-capacity, transformer-based [28, 25] hypernetwork.
-Its added parameter count is $\approx$4.63 M, still less than the most recent competitor [56] and much smaller than [9].
+  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based (Liu et al., 2022) hypernetwork.
+Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like Cao et al.
+(2023).
+  • Thera Pro: The strongest version uses a high-capacity, attention-based hypernetwork.
+Its added parameter count is $\approx$4.63 M, still less than the most recent competitor (Zhu et al., 2025) and much smaller than Chen et al.
+(2023), both attention-based.
 Datasets and metrics.
-Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes.
-We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 [5], Set14 [53], BSDS100 [31], Urban100 [19], and Manga109 [32].
+Following previous work, our models are trained with the DIV2K (Agustsson & Timofte, 2017) training set, consisting of 800 high-resolution RGB images of diverse scenes.
+We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2012), BSDS100 (Martin et al., 2001), Urban100 (Huang et al., 2015), and Manga109 (Matsui et al., 2017).
 Following prior work, we use peak signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric and compute it in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets.
 Additional quantitative results are given in Appendix C.
-Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix F).
+Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix H).
 Backbones.
-Like previous work, we combine our method with two standard backbones for super-resolution and image restoration: (i) EDSR-baseline [26] (1.22 M parameters) and (ii) RDN [54] (22.0 M parameters).
+We combine each of the three variants of our method with two standard backbones for super-resolution and image restoration, as done in previous work: (i) EDSR-baseline (Lim et al., 2017) (1.22 M parameters) and (ii) RDN (Zhang et al., 2018) (22.0 M parameters).
 4.1 Super-Resolution Performance
-Low-res input LIIF [10] SRNO [45] MSIT [56] Thera Pro (ours) GT
-DIV2K
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-DIV2K
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Set14
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Urban100
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Manga109
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN [54] backbone for all methods. Best viewed zoomed in.
-Qualitative results.
-Upon visual inspection – see Fig.
-4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures.
-Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Fig.
-9 in Appendix B).
+Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the held-out DIV2K validation set.
+The highest PSNR value per backbone and scaling factor is bold and the second highest is underlined.
 Backbone (params.) Method Num. of In-distribution Out-of-distribution
@@
 Bicubic 31.01 28.22 26.66 24.82 22.27 21.00 20.19 19.59
-EDSR- baseline [26] (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
-MetaSR [18] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
-LIIF [10] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
-LTE [24] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
-CUF [44] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
-CiaoSR [7] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
-CLIT [9] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
-SRNO [45] 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
-MSIT [56] 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
+EDSR- baseline (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
+MetaSR 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
+LIIF 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
+LTE 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
+CUF 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
+CiaoSR 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
+CLIT 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
+SRNO 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
+MSIT 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
 Thera Air (ours) .008 M 34.75 31.09 29.10 26.84 23.80 22.26 21.26 20.56
@@
 Thera Pro (ours) 4.63 M 35.19 31.50 29.51 27.19 24.09 22.51 21.48 20.73
-RDN [54] (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
-MetaSR [18] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
-LIIF [10] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
-LTE [24] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
-CUF [44] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
-CiaoSR [7] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
-CLIT [9] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
-SRNO [45] 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
-MSIT [56] 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
-Thera Air .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
-Thera Plus 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
-Thera Pro 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
-Table 1: Quantitative comparison of peak signal-to-noise ratio (PSNR, in dB) obtained by various methods on the held-out DIV2K validation set.
-The highest PSNR value per backbone and scaling factor is bold and the second highest is underlined.
-Quantitative results.
-We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above.
-Table 1 shows PSNR values for all tested methods, for both in-distribution (×$\times 2$ to ×$\times 4$) and out-of-distribution (×$\times 6$ to ×$\times 30$) scaling factors.
-Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs.
-29.22 on EDSR ×$\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT [56], and less than a third of CLIT [9].
-Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count.
-This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction.
-For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective ×$\times 4$ backbone.
-This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
+RDN (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
+MetaSR 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
+LIIF 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
+LTE 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
+CUF 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
+CiaoSR 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
+CLIT 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
+SRNO 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
+MSIT 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
+Thera Air (ours) .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
+Thera Plus (ours) 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
+Thera Pro (ours) 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
+Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN (Zhang et al., 2018) backbone.
+The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
 Method Set5 Set14 B100 Urban100 Manga109
@@
 Thera Pro 38.36 34.88 32.79 34.43 30.85 29.08 32.46 29.39 27.87 33.63 29.58 27.26 39.62 34.98 31.98
-Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN [54] backbone.
-The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
-Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Tab.
-2.
+Quantitative results.
+We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above.
+Table 1 shows PSNR values for all tested methods, for both in-distribution ($\times 2$ to $\times 4$) and out-of-distribution ($\times 6$ to $\times 30$) scaling factors.
+Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs.
+29.22 on EDSR $\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT (Zhu et al., 2025), and less than a third of CLIT (Chen et al., 2023).
+Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count.
+This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction.
+For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective $\times 4$ backbone.
+This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
+Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Table 2.
 Our method again outperforms all competing methods in all settings, often substantially (e.g., 29.58 vs.
-29.14 on Urban100 ×$\times 3$).
+29.14 on Urban100 $\times 3$).
 We hypothesize that Thera’s hard-wired PSF is also beneficial when generalizing to unseen datasets.
 Once again we observe that the performance of Thera Air is often comparable to that of methods with orders of magnitude higher parameter counts.
-Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor [10].
+Low-res input LIIF SRNO MSIT Thera Pro (ours) GT
+DIV2K
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+DIV2K
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Set14
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Urban100
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Manga109
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Manga109
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN (Zhang et al., 2018) backbone for all methods. Best viewed zoomed in.
+Qualitative results.
+Upon visual inspection – see Figure 4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures.
+Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Figure 9 in Appendix B).
+Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor (Chen et al., 2021).
 Due to its natural, Fourier-inspired representation, Thera is also infinitely differentiable, while ReLU-based competitors approximate the signal as a piecewise-linear function with null higher derivatives (last row).
 Fidelity of the Signal and its Derivatives.
-Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives [41].
+Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives (Sitzmann et al., 2020b).
 We observe similar effects for Thera, whose thermal activations at $t=0$ can be seen as a special case of periodic activations, cf.
@@
 4.2 Ablation Studies
+Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
 Experiment In-distribution Out-of-distribution
@@
 Thera Plus 34.89 31.22 29.24 26.96 22.34 20.61
-$\times$2 only, fixed $\kappa$ 35.01 31.00 28.84 26.47 21.97 20.33
-$\times$3 only, fixed $\kappa$ 34.74 31.25 29.25 26.91 22.25 20.54
-$\times$4 only, fixed $\kappa$ 34.47 31.17 29.24 26.97 22.32 20.59
-Fixed $\kappa$ 34.86 31.22 29.24 26.95 22.31 20.58
+$\times$2 only, fixed $\kappa$ 35.00 30.96 28.76 26.34 21.87 20.25
+$\times$3 only, fixed $\kappa$ 34.70 31.25 29.25 26.90 22.23 20.53
+$\times$4 only, fixed $\kappa$ 34.36 31.18 29.25 26.98 22.31 20.58
+Fixed $\kappa$ 34.85 31.22 29.24 26.95 22.29 20.56
 No GSE 34.81 31.15 29.18 26.91 22.29 20.57
@@
 Predicted comps. 34.90 31.23 29.25 26.97 22.34 20.61
-Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
 In Table 3, we ablate individual components and design choices of our method to understand their contributions to overall performance.
@@
 $\kappa$ was fixed at the theoretically derived value for these experiments, as multi-scale training is required to optimize it.
-As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Tab.
-3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
+As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Table 3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
 Trainable $\bm{\kappa}$.
-Fixing $\kappa$ at the theoretically derived value (Eq.
-3) leads to a small drop in performance.
+Fixing $\kappa$ at the theoretically derived value (Equation 3) leads to a small drop in performance.
 This suggests that there remain effects that are not accounted for by our proposed observation model, albeit very minor.
 Geometric self-ensemble.
-In line with previous work [43, 26, 44] we see a notable performance boost with geometric self-ensembling.
+In line with previous work (Timofte et al., 2016; Lim et al., 2017; Vasconcelos et al., 2023) we see a notable performance boost with geometric self-ensembling.
 Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
@@
 Thermal activations.
-We replace thermal activations (Eq.
-2) with standard ReLU activations.
+We replace thermal activations (Equation 2) with standard ReLU activations.
 What remains is only the hypernetwork controlling the parameters of the local fields.
@@
 $\text{PSF}(\mathbf{x})=\frac{1}{4\pi\kappa}\cdot\exp\left(-\frac{x_{1}^{2}+x_{2}^{2}}{4\kappa}\right).$ (8)
-In this formulation, we attempt to recover a “pure” signal at $t=0$ or higher sampling rates $10<t<$ 1 given an observation at $t=1$. Note that
-$\Phi(\mathbf{x},t)\text{ is}\begin{cases}\text{meaningless},&\text{if}\ t<0\\ \text{pure signal},&\text{if}\ t=0\\ \text{ill-posed},&\text{if}\ 0<t<1\\ I,&\text{if}\ t=1\\ \text{well-posed},&\text{if}\ t>1\\ \end{cases}$ (9)
-The ill-posed problem for $10<t<$ 1 is the interesting case, where this formulation relates to super-resolution.
+In this formulation, we attempt to recover a “pure” signal at $t=0$ or higher sampling rates $0<t<1$ given an observation at $t=1$. Note that
+[…]
+The ill-posed problem for $0<t<1$ is the interesting case, where this formulation relates to super-resolution.
 The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
@@
 To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$.
-One way to do so is to match the cut-off frequency of the filter in Eq.
-7 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate.
-We take the cut-off frequency of the Gaussian filter defined in Eq.
-7 to be the frequency whose amplitude is halved, which is
+One way to do so is to match the cut-off frequency of the filter in Equation 7 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate.
+We take the cut-off frequency of the Gaussian filter defined in Equation 7 to be the frequency whose amplitude is halved, which is
 $f_{c}=\sqrt{\ln(4)}\cdot\sigma_{f}=\frac{\sqrt{\ln(4)}}{2\pi\sqrt{2kt}}$ (10)
-For the signal compressed into the domain $[-0.5,0.5]$, we can compute the Nyquist frequency to be
+For the signal compressed into the domain [0.5,0.5][-0.5,0.5], we can compute the Nyquist frequency to be
 $f_{\text{Nyquist}}=\frac{N}{2},$ (11)
@@
 If we solve for $f_{c}=f_{\text{Nyquist}}$ at $t=1$, we get
-$\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}N^{2}}.$ (12)
+$\kappa=\frac{\ln(4)}{2\pi^{2}N^{2}}.$ (12)
 For our proposed Thera formulation, we want $\Phi$ to contain a single pixel at $t=1$.
@@
 Therefore we initialize $\kappa$ with
-$\kappa=\frac{\sqrt{\ln(4)}}{2\pi^{2}}.$ (13)
+$\kappa=\frac{\ln(4)}{2\pi^{2}}.$ (13)
 Note that the exact value of $\kappa$ will depend on the characteristics of the system that is being modeled and the anti-aliasing filter that was used (or is assumed).
 Lower values of $\kappa$ allow for sharper signals to be represented at any given value of $t$, but are also more prone to aliasing.
-Finally, we would like to highlight that Eq.
-10 is specific to the case where $\mathbf{x}$ is 2-dimensional.
+Finally, we would like to highlight that Equation 10 is specific to the case where $\mathbf{x}$ is 2-dimensional.
 The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using our field’s formulation in spaces with numbers of spatial dimensions other than 2.
@@
 If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it down to $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$.
-Using Eq.
-10 and Eq.
-11, we can easily find the quadratic relationship
+Using Equation 10 and Equation 11, we can easily find the quadratic relationship
 $t=S^{2}.$ (14)
 For instance, if we want to upsample the image by a factor of 2, we should use $t=0.5^{2}=0.25$.
-Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling.
+Thus, $0<t<1$ refers to super-resolution, while $t>1$ refers to downsampling.
 This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
@@
 Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA
-Neural field
+Heat field
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 6 shows an example where we fit a neural heat field at $t=1$ to the image.
-After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq.
-4.
+After training, any low-pass filtered version of the image can be generated by setting $t$ according to Equation 4.
 We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
@@
 Sampling the center location at $t=0$ would not be representative of the pixel’s footprint.
-Figure 8: Blurring an image before a point-wise sample is equivalent to observing with PSF equivalent to the blur kernel.
-In Fig.
-7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations.
+Figure 8: Blurring an image before a point-wise sample is equivalent to observing with a PSF equivalent to the blur kernel.
+In Figure 7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations.
 Note that aliasing is not always as obvious to the eye as Moiré patterns: in the cases shown in the figure, it would simply mean that sampling the center location at $t=0$ would not be representative of the pixel’s footprint.
@@
 A.5 Initialization of Components
-We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Eq.
-1, is important.
+We have noticed during our experiments that the initialization of the components, $\mathbf{W}_{1}$ in Equation 1, is important.
 Sitzmann et al.
-[41] made similar observations when periodic activation functions were first used for neural fields.
+(2020b) made similar observations when periodic activation functions were first used for neural fields.
 The final distribution of frequencies $\left|\nu(\mathbf{W}_{1})\right|$ did not change much during training.
@@
 Appendix B Continuous Upsampling Example
-In Fig. 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
-Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN [54] backbone, shown with non-integer scaling factors.
+In Figure 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
+Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN (Zhang et al., 2018) backbone, shown with non-integer scaling factors.
 Appendix C Additional Quantitative Results
 C.1 Further Metrics
-In-distribution OOD
-×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$
-LIIF [10] .934 .866 .804 .704 .636
-LTE [24] .936 .868 .807 .707 .639
-CiaoSR [7] .938 .871 .814 .718 .651
-SRNO [45] .937 .873 .819 .738 .685
-MSIT [56] .942 .882 .829 .751 .700
-Thera Pro .943 .884 .833 .757 .706
+Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1.
+We use the SSIM implementation from torchmetrics (Detlefsen et al., 2022).
+We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
 Table 4: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set.
 We use the RDN backbone for all models.
-Some methods did not provide checkpoints at the time of writing, see Appendix F.
+Some methods did not provide checkpoints at the time of writing, see Appendix H.
+In-distribution OOD
+$\times 2$ $\times 3$ $\times 4$ $\times 6$ $\times 8$
+LIIF .934 .866 .804 .704 .636
+LTE .936 .868 .807 .707 .639
+CiaoSR .938 .871 .814 .718 .651
+SRNO .937 .873 .819 .738 .685
+MSIT .942 .882 .829 .751 .700
+Thera Pro .943 .884 .833 .757 .706
+Thera Pro (std.) $\pm$ 2.0e-5 $\pm$ 4.0e-5 $\pm$ 8.9e-5 $\pm$ 9.7e-5 $\pm$ 5.5e-5
+Table 5: Evaluation of GMSD (lower is better) on DIV2K for various methods.
+Backbone Model In-distribution Out-of-distribution
+$\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
+EDSR-b. LIIF 0.0065 0.0354 0.0667 0.1192 0.1976 0.2319 0.2516 0.2645
+LTE 0.0064 0.0351 0.0664 0.1187 0.1970 0.2316 0.2515 0.2645
+Ciao 0.0063 0.0340 0.0641 0.1159 0.1942 0.2295 0.2497 0.2635
+SRNO 0.0063 0.0344 0.0648 0.1165 0.1947 0.2296 0.2499 0.2631
+MSIT 0.0061 0.0334 0.0636 0.1156 0.1944 0.2298 0.2502 0.2634
+Thera Pro 0.0057 0.0314 0.0596 0.1094 0.1878 0.2242 0.2451 0.2591
+Thera Pro (std.) $\pm$ 1.0e-5 $\pm$ 4.1e-5 $\pm$ 7.6e-5 $\pm$ 1.6e-5 $\pm$ 7.1e-5 $\pm$ 5.9e-5 $\pm$ 1.2e-4 $\pm$ 1.6e-4
+RDN LIIF 0.0060 0.0330 0.0628 0.1145 0.1934 0.2285 0.2489 0.2622
+LTE 0.0058 0.0327 0.0622 0.1135 0.1925 0.2280 0.2487 0.2621
+Ciao 0.0058 0.0321 0.0609 0.1114 0.1897 0.2258 0.2470 0.2614
+SRNO 0.0057 0.0321 0.0611 0.1118 0.1906 0.2266 0.2477 0.2613
+MSIT 0.0057 0.0321 0.0610 0.1122 0.1920 0.2288 0.2501 0.2641
+Thera Pro 0.0055 0.0309 0.0588 0.1083 0.1869 0.2235 0.2446 0.2588
+Thera Pro (std.) $\pm$ 7.9e-6 $\pm$ 1.4e-5 $\pm$ 4.0e-5 $\pm$ 3.1e-5 $\pm$ 3.7e-5 $\pm$ 2.0e-5 $\pm$ 6.3e-5 $\pm$ 9.0e-5
+Table 6: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN (Zhang et al., 2018) backbone.
+For some methods, code and/or checkpoints were not publicly available, see Appendix H.
 Method Set5 Set14 B100 Urban100 Manga109
@@
 Thera Pro 29.51 27.34 26.90 25.38 26.12 25.04 24.70 23.24 27.94 25.49
-Table 5: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN [54] backbone.
-For some methods, code and/or checkpoints were not publicly avalibale, see Appendix F.
-Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1.
-We use the SSIM implementation from torchmetrics [13].
-We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
-Furthermore, in Table 5 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
+Table 5 reports Gradient Magnitude Similarity Deviation (Xue et al., 2013) (GMSD) metrics for various methods.
+Thera reaches significantly lower GMSD for all backbone-scale combinations, indicating more faithful gradient structure.
+Note that some methods did not provide code/checkpoints and can therefore not be re-evaluated (see Section H).
+Furthermore, in Table 6 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above $\times 4$).
 C.2 Parameter Efficiency
-In Fig. 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
-$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$34.734.734.734.734.834.834.834.834.934.934.934.93535353535.135.135.135.135.235.235.235.2Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×2 Scaling Factor
-$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$3131313131.131.131.131.131.231.231.231.231.331.331.331.331.431.431.431.431.531.531.531.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×3 Scaling Factor
-$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$2929292929.129.129.129.129.229.229.229.229.329.329.329.329.429.429.429.429.529.529.529.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×4 Scaling Factor
-Figure 10: Comparison of ASR methods for different scaling factors (×2, ×3, and ×4) on DIV2K. Thera consistently achieves better performance at lower parameter counts across all scaling factors.
-Appendix D Hypernetwork Architecture
-In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN [41].
-Thera Plus uses 6 ConvNeXt [29] blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer.
+In Figure 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
+[…]
+[…]
+[…]
+Figure 10: Comparison of ASR methods for different scaling factors ($\times$2, $\times$3, and $\times$4) on DIV2K.
+Thera consistently achieves better performance at lower parameter counts across all scaling factors.
+C.3 Error Bars
+Table 7 reports standard deviations for PSNR observed over N=3N\!=\!3 training runs that were initialized with different random seeds, complementing the main table.
+For SSIM and GMSD, we have reported standard deviations alongside the respective metrics in Tables 4 and 5.
+Standard deviations are in all cases significantly lower than performance improvements over competing methods.
+Table 7: Standard deviation of PSNR (in dB) on the DIV2K validation set over N=3N\!=\!3 runs.
+Backbone Method In-distribution Out-of-distribution
+$\times$2 $\times$3 $\times$4 $\times$6 $\times$12 $\times$18 $\times$24 $\times$30
+EDSR-b. Thera Air 0.0137 0.0143 0.0149 0.0107 0.0084 0.0055 0.0040 0.0040
+Thera Plus 0.0131 0.0122 0.0114 0.0094 0.0075 0.0049 0.0060 0.0022
+Thera Pro 0.0063 0.0088 0.0062 0.0034 0.0054 0.0049 0.0056 0.0034
+RDN Thera Air 0.0187 0.0152 0.0113 0.0045 0.0036 0.0054 0.0022 0.0034
+Thera Plus 0.0045 0.0018 0.0011 0.0011 0.0009 0.0013 0.0026 0.0024
+Thera Pro 0.0025 0.0009 0.0038 0.0024 0.0021 0.0036 0.0021 0.0012
+Appendix D Analysis of Learned Components & Kappa
+Figure 11 shows a statistical analysis of the frequency components learned by Thera-Pro with an RDN backbone for the DIV2K training data.
+Components are distributed uniformly across directions, with fewer components at low frequencies and progressively more components at higher frequencies.
+This provides more representational capacity for the reconstruction of fine details, analogous to the typical distribution of frequency components for other decompositions such as the Fourier or discrete cosine transforms.
+To investigate whether these components form a generalizable basis, we fit heat fields to images from various datasets other than the training data: Urban100, Manga109, and a highly out-of-distribution subset from the fastMRI (Zbontar et al., 2018) medical imaging dataset (single-coil knee validation split, with intensities scaled to [0,1]).
+In all experiments, we fix the frequency bank to the one shown in Figure 11 and optimize only the scale and shift parameters with the AdamW optimizer for 600 iterations per image, with learning rate 0.001.
+We use local fields that cover patches of $N\times N$ pixels, with $N\in\{3,4,6,12,18\}$.
+Table 8 shows that the pre-trained components reconstruct all datasets with negligible error at smaller local field sizes (in-distribution scales), and still achieve very low errors at larger field sizes up to $18\times 18$ pixels (out-of-distribution scales).
+These numbers are not surprising, as the frequency bank was optimized to fit local fields and can act as an over-parametrized dictionary for smaller field sizes (e.g., $4\times 4$ GT pixels per field).
+Notably, it still works relatively well for large OOD sampling factors.
+Also, there is no obvious difference between the values for natural images and those for MRI images, which suggests that the frequency bank is a general-purpose basis, similar to Fourier or DCT components.
+Table 8: Reconstruction MSE when fitting heat field grids of various sizes to out-of-distribution datasets.
+Local field size $\times3$ $\times4$ $\times6$ $\times12$ $\times18$
+Set5 2e-10 6e-11 1e-11 1.37e-6 2.16e-4
+Urban100 4e-9 9e-10 2e-10 2.83e-6 7.54e-4
+Manga109 2e-8 4e-9 7e-10 1.41e-6 3.01e-4
+fastMRI 3e-11 1e-11 1e-11 8.83e-6 9.23e-4
+Furthermore, Table 9 reports final, converged values of the thermal diffusivity coefficient $\kappa$.
+We observe that $\kappa$ is very similar across runs ($\sigma\approx 0.003$), but deviates from the theoretically derived value (for a Gaussian downsampling model, Equation 3) by a factor of $\approx 1/2$.
+This deviation suggests that the cutoff frequency of the anti-aliasing used in the cubic Mitchell-Netravali filter – used to downsample images during training – does not exactly match the Gaussian one used in the theoretical derivations, being more lenient towards aliasing than the theoretical model (anti-aliasing filters are tuned empirically to balance aliasing against loss of details).
+The result indicates that Thera is indeed able to tune $\kappa$ as necessary to approximate the downsampling characteristics seen in the training data, in a repeatable manner.
+Table 9: Converged values of $\kappa$ for multiple runs.
+Backbone Model Converged $\kappa$
+EDSR-b. Thera Air 0.0295
+Thera Plus 0.0336
+Thera Pro 0.0266
+RDN Thera Air 0.0300
+Thera Plus 0.0360
+Thera Pro 0.0293
+Theoretical Gaussian 0.0702
+Figure 11: Statistical distribution of the converged frequency bank of the Thera Pro run with an RDN backbone.
+Left: Polar scatter plot of spatial frequency and angular direction of components.
+Right: Corresponding marginal distributions.
+Appendix E Hypernetwork Architecture
+In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN (Sitzmann et al., 2020b).
+Thera Plus uses 6 ConvNeXt (Liu et al., 2022) blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer.
 Projection blocks are added between blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
-For Thera Pro, two SwinIR [25] blocks are used with depth 7 and 6, respectively, and 6 attention heads per layer.
+For Thera Pro, two windowed transformer-based blocks (Liang et al., 2021; Liu et al., 2021) blocks are used with 7 and 6 layers respectively, and 6 attention heads in each layer.
 There are 128 field parameters for Thera Air and 2048 for the larger variants Thera Plus and Thera Pro.
@@
 We highlight that these do not come from the hypernetwork.
-Appendix E Real-World Optical Zoom Data
-To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset [14].
-We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) [21] that estimates just 6 parameters per image to correct domain-specific affine distortions.
+Appendix F Computational Complexity
+Table 10 reports inference time and VRAM requirements of different Thera variants as well as comparison methods, on the $48\times 48$ pixels standard input patch size and scaling factor 4 (i.e., $192\times 192$ output size).
+Each Thera variant improves compute time and memory efficiency compared to methods with similar or higher parameter counts, with particularly pronounced gains over Transformer-based competitors (e.g., Thera Pro uses less than $1/4$ of the time and $1/13$ of the VRAM footprint of MSIT, at similar parameter count).
+All tests were performed on an NVIDIA GeForce RTX 3090 Ti GPU.
+Table 10: Comparison of runtime and VRAM footprint of various methods with an EDSR-baseline backbone.
+Method Add. params. Time VRAM
+MetaSR 0.45 M 3.72 ms 834 MiB
+LIIF 0.35 M 13.8 ms 604 MiB
+LTE 0.49 M 16.7 ms 574 MiB
+CiaoSR 1.43 M 37.5 ms 3894 MiB
+CLIT 15.7 M 86.5 ms 9402 MiB
+SRNO 0.80 M 11.2 ms 2702 MiB
+MSIT 4.83 M 92.2 ms 9388 MiB
+Thera Air .008 M 3.56 ms 322 MiB
+Thera Plus 1.41 M 14.2 ms 664 MiB
+Thera Pro 4.63 M 19.44 ms 686 MiB
+Appendix G Real-World Optical Zoom Data
+Table 11: Results (PSNR in dB) on the COZ test set.
+Method In-distribution Out-of-distribution
+$\times$2 $\times$3 $\times$4 $\times$5 $\times$6
+MetaSR 28.70 26.55 25.17 24.31 23.25
+LIIF 28.72 26.61 25.16 24.32 23.23
+LTE 28.67 26.55 25.15 24.37 23.26
+SRNO 28.73 26.59 25.15 24.31 23.25
+LIT 28.74 26.58 25.15 24.35 23.19
+LMI 28.86 26.66 25.22 24.39 23.29
+Thera++ 29.06 26.84 25.45 24.49 23.46
+To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset (Fu et al., 2024).
+We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) (Jaderberg et al., 2015) that estimates just 6 parameters per image to correct domain-specific affine distortions.
 Despite its name, the STN is not a transformer network in the modern sense that employs attention mechanisms; rather, it is an image model block that explicitly allows the spatial manipulation of data within a convolutional neural network.
@@
 These challenges make super-resolution particularly difficult for this dataset.
-However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Tab.
-6.
-Method In-distribution Out-of-distribution
-$\times$2 $\times$3 $\times$4 $\times$5 $\times$6
-MetaSR [18] 28.70 26.55 25.17 24.31 23.25
-LIIF [10] 28.72 26.61 25.16 24.32 23.23
-LTE [24] 28.67 26.55 25.15 24.37 23.26
-SRNO [45] 28.73 26.59 25.15 24.31 23.25
-LIT [9] 28.74 26.58 25.15 24.35 23.19
-LMI [14] 28.86 26.66 25.22 24.39 23.29
-Thera++ 29.06 26.84 25.45 24.49 23.46
-Table 6: Results (PSNR in dB) on the COZ test set.
-Appendix F Reproducibility of Existing Methods
+However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Table 11.
+Appendix H Reproducibility of Existing Methods
 We encountered challenges attempting to recreate the results reported for some of the competing methods, which explains why some are missing or differ from the originally reported numbers.
 Details are provided below.
-CUF [44].
+CUF (Vasconcelos et al., 2023).
 At the time of writing, there were no public code or checkpoints available for CUF.
@@
 Furthermore, we could not create any qualitative samples using CUF.
-CLIT [9].
+CLIT (Chen et al., 2023).
 For CLIT, at the time of writing, there is a public code, but no checkpoints.
-We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using 8×8$\times$8 × Nvidia GeForce RTX 3090 GPUs.
+We have made a bona fide attempt to reproduce the models, but due to the cascaded training schedule and the large model size, the training process would require excessive amounts of compute: over a month using $8\times$ Nvidia GeForce RTX 3090 GPUs.
 Unfortunately, the authors did not respond to our requests for the trained checkpoints used in their paper.
-CiaoSR [7].
+CiaoSR (Cao et al., 2023).
 We found that in the official CiaoSR implementation border cropping prior to evaluating on DIV2K deviated slightly from all other methods.
@@
 All DIV2K numbers were re-computed with the corrected code, resulting in slight deviations from those reported in the original paper.
-MSIT [56].
-The numbers reported in the paper were achieved by training on a roughly 3$\times$ larger training set, comprising not only DIV2K but also Flickr2K.
-For a fair comparison with all other ASR methods used in our paper, we re-trained MSIT on DIV2K alone.
-We used the authors’ official code and configuration files and trained for two stages, each comprising 1050 epochs.
-For the second (RiM) training stage we used scaling factors between ×$\times 1$ and ×$\times 4$.
-Performance generally drops as a result of the smaller training set, in line with the ablation experiments reported for MSIT, where the models were also trained only with DIV2K.
+[…]